### v1

# Instance-level Performance Prediction for Long-form Generation Tasks

Abstract
We motivate and share a new benchmark [^1] Benchmark URL forthcoming for instance-level performance prediction of long-form generation tasks having multi-faceted, fine-grained quality metrics. Our task-, model- and metric-agnostic formulation predicts continuous evaluation metric scores given only black-box model inputs and outputs. Beyond predicting point estimates of metric scores, the benchmark also requires inferring prediction intervals to quantify uncertainty around point estimates. Evaluation spans 11 long-form datasets/tasks with multiple LLMs, baselines, and metrics per task. We show that scores can be effectively predicted across long-form generation tasks using as few as 16 training examples Overall, we introduce a novel and useful task, a valuable benchmark to drive progress, and baselines ready for practical adoption today.
**footnotetext: Both authors contributed equally.****footnotetext: Work does not relate to the author’s position at Amazon.
1 Introduction
Large language models (LLMs) can tackle increasingly challenging open-ended, long-form generation tasks across diverse modalities (e.g., text, images, code, etc.) Liu et al. (2024); Yin et al. (2024). Given the vast output space, many different generations of varying quality are possible Giulianelli et al. (2023). This motivates evaluation of long-form output quality that is: 1) fine-grained and nuanced (rather than binary correct/incorrect), and 2) multi-faceted, with diverse evaluation metrics used to assess different aspects of output quality Gao et al. (2025). For example, summarization quality might be measured via aspects such as brevity, fluency, coherence, consistency, and relevancy.
Like other AI models, LLMs are fallible, with varying output quality across different inputs Schellaert et al. (2025). Consequently, the ability to predict the quality of each output – across different inputs, tasks, models, and metrics – would have myriad benefits. Key among these is to guide appropriate trust and use for both human users Steyvers et al. (2025); Dhuliawala et al. (2023); Bhatt et al. (2021) and downstream agents Mielke et al. (2022); Yoffe et al. (2024). For example, outputs predicted to be of high quality might be used as-is, while those of lesser quality might be revised or discarded accordingly. Predicted quality could also provide a signal to optimize prompt selection or model parameterization Zhou et al. (2022), to select between different generations Chen et al. (2023), or to route inputs to whichever model is predicted to perform best Lee et al. (2024); Šakota et al. (2024); Lu et al. (2023); Shnitzer et al. (2023); Ding et al. .
Assuming evaluation metrics for fine-grained output quality are continuous $\in[0,1]$, we not only want to predict point estimates for metric scores, but also prediction intervals to quantify uncertainty around point estimates Shrestha and Solomatine (2006); Patel (1989); Geisser (1993). For example, given a point estimate prediction of 0.8, a tight 95% interval (say [0.79, 0.81]) would predict that output quality is quite close to the point estimate, while a loose 95% interval (say [0.6, 0.1]) would mean that quality may actually be much lower. Prediction intervals thus provide crucial uncertainty quantification for risk-averse settings by indicating probabilistic lower bounds on expected performance Harrington et al. (2019). We thus emphasize the importance of prediction intervals as well as point estimates for performance prediction.
In this work, we motivate and formulate the task of instance-level performance prediction for long-form generation tasks: predicting instance-level output quality – point estimates and prediction intervals – for continuous evaluation metrics. By design, our task formulation is task-, model-, and metric-agnostic: given an input-output pair for any black-box model, one must predict arbitrary evaluation metrics. The generality of this formulation supports broad use across the diverse, ever-changing landscape of LLM models and evaluation metrics for long-form output quality. We assume supervision is key: by observing examples of past input-output pairs and corresponding metric scores, one can learn a metric-specific prediction function. An ideal solution would achieve strong predictive performance with minimal supervision.
To promote research on instance-level performance prediction for long-form generation tasks, we create and share a new benchmark spanning twelve long-form tasks/datasets: text summarization, machine translation, code generation, and question answering. With different metrics often used to assess different aspects of output quality, our benchmark includes multiple evaluation metrics for most tasks. In regard to LLM coverage , we include outputs from Llama 3.2 ;other LLMs could be easily added later.
Our benchmark includes two baselines. One uses off-the-shelf confidence estimators (verbalized confidence and output consistency) to elicit weak features, then trains a classic regression model (e.g., linear regression, random forests) to predict the metric score from the features. The other baseline uses a reference-free LLM-as-a-Judge Li et al. (2024a) with in-context learning for textual regression Vacareanu et al. (2024) to directly predict the metric score from the input-output pair.
Our work demonstrates the feasibility of instance-level prediction of continuous evaluation metric scores and associated uncertainty intervals. We show that scores can be predicted with low RMSE and CRPS, across long-form generation tasks, and using as few as 16 training examples (Section 5.2). While the feature-based regression baseline described above is shown to perform well, the LLM judge baseline proves inconsistent. We also see that some confidence-based features are much more predictive than others. Taken together, our formulation, benchmark, and findings provide both a strong foundation for future research as well as baselines suitable for practical use today.
2 Related Work
Our formulation of performance prediction draws upon a large body of prior work across related areas: Confidence Estimation and Uncertainty Quantification (Section 2.1), Score and Benchmark Predictability (Section 2.2), Query Performance Prediction in information retrieval (Section 2.3), Quality & Metric Estimation in machine translation (Section 2.4), Textual Regression (Section  2.5), and LLM-as-a-judge & Reward Models (Section 2.6).
2.1 Uncertainty and Confidence Estimation
Our work is particularly motivated by confidence estimation (CE) Xiong et al. (2024) and uncertainty quantification (UQ) Vashurin et al. (2025); Xiao et al. (2022), which similarly predict how much each system output can be trusted. While terminology and definitions vary across different works, UQ (Smith, 2013) canonically seeks to quantify certainty regarding a given process, observation, or system output, while CE seeks to predict the probability that a system output is correct or not (Guo et al., 2017; Steyvers et al., 2025).
However, as Huang et al. (2024) write, “the correctness of long-form generation is not always either true or false but can be partially correct. Therefore, a single confidence score for long-form outputs is ambiguous: it can either imply ‘I am x% confident that the answer is 100% correct’ or ’I am 100% confident that the answer is x% correct.” Moreover, Ulmer et al. (2024) suggest estimation of LLM confidence intervals as future work.
Nevertheless, prior work has typically modeled long-form generation confidence via binary correctness by: i) mapping generations to a most similar multiple choice answer (Tjuatja et al., 2024), ii) binarizing partial-credit scores via thresholding (Kuhn et al., 2023; Lin et al., 2024; Band et al., 2024; Pedapati et al., 2024), iii) testing whether or not the output contains an answer (string) (Rajpurkar et al., 2016; Yu et al., 2018), iv) focusing on tasks with short answers wherein evaluation can rely on exact matching (Pedapati et al., 2024; Xiao et al., 2022); or v) decomposing a long output into smaller nuggets checked individually for correctness (Zhang et al., 2025, 2024a; Min et al., 2023).
In recent work, Zhang et al. (2025) usefully infer partial-credit scores for correctness at various granularities (e.g. sentence or passage level) but still assume that output quality can be measured via a single dimension of correctness. Huang et al. (2024) also emphasize “…the multifaceted nature of evaluating long-form outputs, where factors like factuality, coherence, clarity, and comprehensiveness each play a role…” However, they assume evaluation dimensions are implicit and latent, leading to subjective disagreements between evaluators. In contrast, we assume distinct evaluation metrics assess different aspects of output quality, and predict each separately. Prior studies also tend to show only correlation between CE vs. evaluation metric scores, while we predict metric scores directly.
Our formulation pursues uncertainty quantification via prediction intervals that quantify uncertainty around point estimates of metric scores.
2.2 Score & Benchmark Predictability
Like CE and UQ above, score predictability work Hernández-Orallo et al. (2022); Pacchiardi et al. (2025); Schellaert et al. (2025) similarly makes instance-level predictions of binary output correctness. Hernández-Orallo et al. note continuous prediction is more complex and proceed with binary. Schellaert et al.  suggest future non-binary work while Pacchiardi et al. assume the binary case.
Three other differences are noteworthy. First, score predictability emphasizes anticipatory prediction of instance-level success before output generation. This avoids computing outputs but makes prediction more difficult (without need for our use cases). Our performance prediction task more closely resembles UQ and CE in which confidence is inferred for a given output. Second, model details are often exploited to improve prediction, vs. our model-agnostic, black-box formulation. Third, score predictability methods tend to leverage a large feature set (e.g. text embeddings); we require only (input, output, score) triples.
Other work on benchmark prediction infers continuous scores, but only for entire datasets in aggregate Ye et al. (2023); Zhang et al. (2024b). Instead, we predict the quality of individual outputs.
2.3 QPP in Information Retrieval
Our nomenclature of “performance prediction” is inspired by query performance prediction (QPP) work in information retrieval (Cronen-Townsend et al., 2002; He and Ounis, 2006; Arabzadeh et al., 2024). Given an input query, QPP seeks to predict the quality of output search results. Like our work, QPP is both instance-based and predicts arbitrary, continuous-valued metrics of output quality.
QPP work is split into pre-retrieval vs. post-retrieval methods. Like anticipatory score predictability (Section 2.2), pre-retrieval methods require predicting output quality (of search results) before executing the (input) query. In contrast, post-retrieval methods run the query and analyze output search results to better predict their quality, and this helps them to perform much better in practice. We also expect output analysis to help prediction.
2.4 Quality and Metric Estimation in MT
Quality estimation (QE) in machine translation (MT) predicts translation output quality at run-time Specia et al. (2009, 2010). Earlier work on CE for MT also assumes binary correctness Blatz et al. (2004) which QE then generalizes to continuous, partial-credit evaluation. The term CE is also still sometimes used to refer to the broader QE task Chelba et al. (2020). QE is canonically framed as a regression task to predict human quality ratings (i.e., “direct assessment”) of output quality.
Scarton et al. (2016) frame QE more broadly as predicting output quality across diverse NLP tasks, including text summarization and simplification as well as QE. Some QE work has also emphasized the value of inferring prediction intervals beyond simple point estimates Glushkova et al. (2021); Giovannotti (2023); Zerva and Martins (2024).
Zouhar et al. (2023) propose “metric estimation”: predicting evaluation metric scores rather than human quality ratings. However, their goal in doing so is to pre-train a QE model. We are agnostic as to the type of prediction target (i.e., metric score or human rating) used to evaluate long-form outputs. To illustrate this, we also include a “direct assessment” task in our evaluation (Section 5.1), complementing our other metric prediction results.
2.5 Textual Regression
Wang et al. (2022) note “almost no work in NLP on calibration in a regression setting” and consider three regression tasks: semantic similarity, sentiment analysis, and QE. They use BERT/SBERT and report point estimates and prediction intervals.
LLM-based work has pursue three primary approaches: i) extracting LLM features for classic regression Tang et al. (2025a), ii) leveraging in-context learning Vacareanu et al. (2024); and iii) fine-tuning for specific regression tasks Tang et al. (2025b). Vacareanu et al. (2024) consider three synthetic regression tasks and show effective results across several models via in-context learning, but without prediction intervals. We report a similar in-context LLM baseline in this work (Section 4.3). Most recently, Vedula et al. (2025) propose quantile regression for LLM-based price estimation, reporting prediction intervals as well as point estimates.
2.6 LLM-as-a-Judge & Reward Models
The rise of LLMs has also brought the advent of LLM-as-a-judge. A key distinction exists between reference-free (RF) methods, suitable for predicting output quality at run-time, vs. reference-based (RB) methods that require human reference outputs to compare AI outputs against Sheng et al. (2024); Li et al. (2024a). Tan et al. (2024) present an excellent discussion relating RF methods to reward models used in reinforcement learning from human feedback (RLHF), noting “these two fields are closely related and are gradually converging.” Our work is related to both fields, though we focus on run-time prediction of RB metrics.
We also note similarity between RF LLM judges and “verbalized” CE methods (Section 2.1) in which an LLM is prompted for its confidence that a given output is correct (Tian et al., 2023; Xiong et al., 2024). Both support run-time use of LLMs through prompting, though verbalized CE methods have assumed a particular type of assessment (i.e., confidence) while RF LLM judges have been conceived and pursued much more broadly.
While RF LLM judges are increasingly being used as final evaluators, their reliability remains contested: they can exhibit self-preference and positional biases Wang et al. (2024); Li et al. (2024b), mis-rank instances they cannot themselves solve without a human reference answer Krumdick et al. (2025), and underperform or distort correctness in domains like software engineering Zhou et al. (2025). More broadly, RF methods have inherent structural biases—being effectively one generation model evaluating another—which risks test-time gaming, favoring models similar to the judge, and even disadvantaging higher-quality (including human) outputs Deutsch et al. (2021). Reproducibility can also be limited by underlying changes in closed LLM judges or prompt sensitivity in use of open judges Bavaresco et al. (2025).
Meta-evaluation of RF methods often reports correlation with human judgments, but caution is warranted. For example, the breadth and diversity of correlation studies is limited Bavaresco et al. (2025). Krumdick et al. (2025) note correlation can also mask disagreement on difficult instances, limiting ability to evaluate more advanced models or reliably predict instance-level quality. Dorner et al. (2025) show that when a RF LLM judge has comparable accuracy to an evaluated model, “no debiasing method can decrease the required amount of ground truth labels by more than half… [showing] severe limitations of the [RF] LLM-as-a-judge paradigm…" Their empirical results show practical performance is even worse than this theoretical limit. Gera et al. (2025) similarly report, “Crucially, even very good instance-level capabilities do not guarantee accurate model ranking.”
While we expect RF LLM judges to both continue to improve and become better understood, the limitations above suggest that RB LLM judges (and perhaps classic RB metrics) still remain an important part of the NLP evaluation toolkit for evaluating long-form outputs. This motivates a continuing need for run-time prediction of RB evaluation metrics, as we pursue in this study.
3 Task Formulation
Prediction intervals represent a natural extension of uncertainty for point estimates, bounding the range within which the true performance score will fall with a specified probability (e.g. 95 %) Shrestha and Solomatine (2006); Patel (1989); Geisser (1993). Prediction intervals are increasingly being reported, e.g., in MT quality estimation (Section 2.4) and other textual regression tasks (Section 2.5).
Formalization. Let $\mathcal{X}$ denote the input space (user queries, upstream agent inputs) and $\mathcal{Y}$ the long-form output space (e.g., text, code, etc.) for both LLM generations and gold reference outputs. We assume each task $t\in\mathcal{T}$ has $K_{t}$ continuous, reference-based (RB) evaluation metrics $\mathcal{M}_{t}=\{m_{t,1},\dots,m_{t,K_{t}}\}$, where each metric $m_{t,k}:\mathcal{Y}\times\mathcal{G}\rightarrow\mathbb{R}$ assigns LLM output $y$ a real-valued quality score based on some measure of agreement with reference gold output $g$. At inference, only input $x$ and generation $y$ are observed; gold output $g$ and metric score $m_{t,k}(y,g)$ are not.
Instance-level performance prediction. For each task-specific metric $m_{t,k}$, we seek a predictor
$f_{t,k}:(\mathcal{X}\!\times\!\mathcal{Y})\rightarrow\big{(}\hat{s},[\ell,u]\big{)}$
for instance $(x,y)$ that returns: (i) a point estimate $\hat{s}\in\mathbb{R}$ for metric score $m_{t,k}(y,g)$ and (ii) a $(1-\alpha)$ prediction interval $\ell,u]\subseteq\mathbb{R}$ around $\hat{s}$ such that:
$\Pr\!\big{(}m_{t,k}(y,g)\in[\ell,u]\big{)}\geq 1-\alpha,$
for a specified confidence level (e.g., $\alpha=0.05$).
Supervision. For each task $t$, assume we have a dataset $\mathcal{D}$ containing $N_{t}$ training instances $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N_{t}}$ for each of $t$’s $K_{t}$ metrics, where $s_{i}=m_{t,k}(y_{i},g_{i})$ denotes score $s_{i}$ for metric $m_{t,k}$, as calculated for generation $y_{i}$ with respect to gold reference $g_{i}$. This setup is both metric-agnostic (allows any metric) and model-agnostic (makes no use of model parameters or decoding internals).
Objective: minimize point-estimation error and also produce calibrated intervals (see Section 5.2).
Figure 1: Illustrating different regression model predictions (solid lines) and confidence intervals (color shading around the lines). In this example, a Bayesian Ridge regression model performs best as predictions are neither over-confident (XGBoost) nor under-confident (Gaussian Process).
4 Baseline Methods
Having defined the desired performance predictor $f_{t,k}$ (Section 3), we now describe two baseline methods for instantiating such a predictor.
Confidence-based regression (CE-Reg): Extract unsupervised confidence estimates as features, then learn a regression model to map features to the continuous metric score and a prediction interval. Methods are presented in Sections 4.1 and 4.2.
Reference-Free LLM-as-a-Judge: Prompt an LLM to directly output a point estimate and prediction interval for the target metric, using labeled examples for in-context learning. See Section 4.3.
4.1 Confidence Estimation (CE)
Verbalized confidence (1S and 2S): Prompting an LLM for its confidence that an output is correct (see Section 2.6). One-step (1S) produces both output and estimated confidence in one invocation, while two-step (2S) first generates the output and then calls the LLM a second time to estimate output confidence. While 1S reduces computation, 2S provides cleaner delineation for benchmarking generation vs. confidence estimation. Note that while 2S permits different LLMs to be used for each step, our experiments use the same LLM for both steps.
Consistency-based: We generate $n$ outputs and measure coherence between them via 5 methods: Degree Matrix (Deg; Lin et al. (2024)), Eccentricity (Ecc; Lin et al. (2024)), Sum of Eigenvalues of the Graph Laplacian (EigV; Lin et al. (2024)), Lexical Similarity (LexSim; Fomicheva et al. (2020)), and the Number of Semantic Sets (NumSet; Lin et al. (2024)). Most use natural language inference (NLI) models to assess coherence between outputs, while LexSim is computed using ROUGE-L.
4.2 Confidence-Based Regression (CE-Reg)
Given a dataset $\mathcal{D}=\{(c_{i},s_{i})\}_{i=1}^{N},$ of $N$ training examples, where each $c_{i}$ is the confidence feature (verbalized or consistency-based) for output $y_{i}$ and $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold score, we fit an auxiliary regression model to learn the conditional distribution $s_{i}\sim p(s\mid c_{i},\theta),$ where $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold metric score. From the fitted distribution we extract a point estimate $\hat{s}_{i}$ and a $(1-\alpha)$ prediction interval $\ell_{i},u_{i}\,]$ (using $\alpha=0.05$) via
$\ell_{i}=F^{-1}\bigl{(}\tfrac{\alpha}{2}\bigr{)},\quad u_{i}=F^{-1}\bigl{(}1-\tfrac{\alpha}{2}\bigr{)}$
where $F$ is the model’s predictive cumulative distribution function.
Regression Models. We compare a variety of models: linear (Linear, Bayesian Ridge), tree ensembles (Random Forest, XGBoost), and a Beta regression for bounded metrics. Gaussian-error models provide mean $\hat{s} &$ variance $\hat{\sigma}^{2}$ directly; ensembles approximate variance via bootstraps/trees.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.45 0.43 0.44 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.45 0.45 0.44 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.15 0.14 0.16 LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.25 0.22 0.23 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.24 0.21 0.21 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.08 0.08 0.08 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.31 0.32 0.34 Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.35 0.34 0.35 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.27 0.28 0.27 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.17 0.16 0.17 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.22 0.22 0.20 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.18 0.17 0.17 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.09 0.09 0.10 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.33 0.32 0.31 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.23 0.22 0.22 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.19 0.18 0.18 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.09 0.09 0.09 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.20 0.20 0.20 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.12 0.12 0.12 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.22 0.20 0.19 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.22 0.20 0.19 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.32 0.34 0.36 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.21 0.20 0.20 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.06 0.06 0.06 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.39 0.40 0.38 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.28 0.29 0.27 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.15 0.12 0.12 AVERAGE 0.14 0.15 0.15 0.14 0.16 0.15 0.15 0.23 0.22 0.22
Table 1: RMSE Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings. BiGGen-BR = BiGGen-Bench-Results. For RMSE, values lower the better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.26 0.25 0.26 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.33 0.34 0.32 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.08 0.08 0.08 ASQA LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.14 0.14 0.14 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.14 0.13 0.14 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.03 0.03 0.03 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.20 0.20 0.22 BigCodeBench Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.23 0.23 0.23 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.18 0.19 0.18 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.10 0.10 0.10 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.14 0.14 0.13 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.11 0.10 0.10 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.03 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.23 0.22 0.20 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.14 0.14 0.13 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.12 0.11 0.11 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 Opus-100 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.11 0.10 0.10 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.07 0.07 0.07 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.13 0.12 0.11 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.27 0.29 0.31 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.27 0.28 0.26 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.19 0.20 0.18 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.06 0.05 0.05 AVERAGE 0.07 0.11 0.08 0.07 0.12 0.12 0.12 0.14 0.14 0.14
Table 2: CRPS Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings. For CRPS, lower is better.
4.3 Reference-free LLM-as-a-Judge
Building on prior work (Section 2.6), we define a Reference-free LLM-as-a-Judge (RF-LLMaaJ) baseline for continuous-score performance prediction using in-context learning Vacareanu et al. (2024). It is prompted with: 1) a brief description of the task and the scoring metric; 2) $N$ demonstration examples in form (input, output, gold metric score) of the desired scoring format; 3) instructions to produce a point score and 95% upper/lower interval bound; and 4) A test pair (input, output) for which it must predict a point estimate and prediction interval for the target metric.
5 Benchmark
5.1 Experimental Setup
Datasets & Tasks. Table 5 details our datasets, tasks, and evaluation metrics used. Tasks span code generation, fact-checking, function-calling, machine translation (MT), question-answering (QA), text summarization, and others. We randomly sample up to 5,000 instances per dataset, splitting each 3/1/1 into train, dev, and test sets. For each question, the LLM generates three outputs. Each output is evaluated based on task-specific metrics, producing a ⟨input, LLM output, metric score⟩ tuple.
Task-Specific Evaluation Metrics. Each task requires predicting one or more continuous metrics. Given our metric-agnostic formulation, the choice of metrics used is less important than assessing the ability to predict arbitrary metrics, including different metrics for the same generation task.
Metrics include: BERTScore (Zhang* et al., 2020), CodeBLEU (Ren et al., 2020), and reference-based (RB) LLM-as-a-judge metrics, such as LLM-Eval (accuracy and informativeness) for QA (Xian et al., 2025; Lin and Chen, 2023) and BiGGen (Kim et al., 2025), GEMBA_ref (with references) (Kocmi and Federmann, 2023) for MT, G-Eval (coherence, consistency, fluency, and relevancy) (Liu et al., 2023) for text summarization, FActscore (Min et al., 2023) for fact-checking, and CodeJudge Tong and Zhang (2024) for code generation and function-calling (functional correctness and inconsistency). These LLM judge methods support more robust evaluation by considering diverse aspects of quality (e.g., factual correctness, informativeness, fluency, relevance, etc.).
We also include a “direct assessment” (DA) task (predicting human-annotated quality ratings instead of automatic metric scores, see Section 2.4), for the BiGGen-Bench-Results dataset (Kim et al., 2025). Because this dataset provides human evaluations for its pre-generated LLM outputs, we use both these annotations and outputs for this DA task. However, because the dataset only contain one output per rating, this precludes evaluation of consistency-based methods (see Section 4.1).
LLMs. Llama 3.2 11B (Meta, 2024) is used in both generating outputs and in our baseline predictors (i.e., for consistency and verbalized confidence methods, as well as our RF-LLMaaJ). We set max tokens =769=769, temperature =1=1, and topP =0.9=0.9. To avoid risk of self-enhancement bias (Zheng et al., 2023) in evaluation, we use a different LLM, Gemini-2.0-Flash-Lite, for reference-based (RB) LLM-as-a-judge evaluation metrics.
Our proof-of-concept direct assessment evaluation setting differs by using human quality ratings and existing LLM outputs: Kim et al. (2025)’s GPT-4-0125 (Achiam et al., 2023) generations.
Baselines. Consistency methods are implemented by LM-Polygraph (Fadeeva et al., 2023) and computed over n=3n\!=\!3 outputs, with all $n$ outputs sharing the same consistency score. We implement regression models in scikit-learn and use XGBoost with default hyperparameters. Beta regression is implemented in-house. For regression model selection, we compare RMSE and CRPS results on the dev set. Based on this (Figure 3), we select Random Forest for our test set evaluation.
Regarding scale of supervision, for RF-LLMaaJ in-context learning, we evaluate varying $N\in\{4,8,16\}$ demonstration examples. We only provide up to 16 examples to the RF-LLMaaJ prompt due to high token costs (especially with long-form outputs), and potentially even degrade performance Levy et al. (2024); Modarressi et al. (2025). In contrast, CE-R is trivial to train at scale.
As for relative LLM compute costs of different methods (beyond the output generation call): consistency methods require n1n-1 additional calls (to generate additional outputs for measuring output consistency), verbalized 1S requires no additional calls, verbalized 2S requires one additional call, and RF-LLMaaJ also requires one additional call. In practice, the $n$ outputs generated for consistency would be used to select the output predicted to perform best, so the added compute cost is offset by expected quality improvement in output selection.
Appendix A.2 lists prompts and parsing rules.
5.2 Evaluating Performance Prediction
$f_{t,k}$ (Sec. 3) must return both a point estimate $\hat{s}$ and a predictive distribution (from which we take a $(1-\alpha)$ interval), and we evaluate these separately.
Point estimates. We report RMSE between prediction $\hat{s}$ and gold score $q$, providing an interpretable measure of average error magnitude in the target’s original units and serving as our primary evaluation metric. Pearson Correlation (PC) is shown only as a diagnostic for feature usefulness (Appendix 9); it does not directly reflect absolute error, and correlation may be high while still incurring large systematic errors. However, as it is unscaled, it is useful for comparing between tasks and metrics.
Prediction intervals. Expected Calibration Error (ECE) is popular for classification, measuring how well predicted probabilities match observed frequencies via binning (Naeini et al., 2015). However, ECE is designed for binary scores, not for continuous ones. Prediction intervals for continuous scores can be visualized in Figure 1, showing each instance’s true score on the y-axis as well as the prediction regressor on the x-axis. Ideally, predictions are not too narrow (overconfident) or too wide (underconfident), and several metrics exist for assessing such miscalibrations.
We use the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007) as our primary metric for evaluating prediction intervals. It is a proper scoring rule that jointly rewards calibration and sharpness of the full predictive CDF $F$, where lower CRPS is better:
$\mathrm{CRPS}(F,s)=\int_{-\infty}^{\infty}\big{(}F(x)-\mathbf{1}\{x\geq q\}\big{)}^{2}\,dx$
For interpretability, we also measure Average Coverage Error (ACE) (Shrivastava et al., 2016; Xie et al., 2023), which compares empirical coverage of the nominal $(1-\alpha)$ interval to its target as $\mathrm{ACE}=\bigl{|}\mathrm{PICP}(p)-p\bigr{|}$ where
$\mathrm{PICP}(p)$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\,\ell_{i}(p)\leq s_{i}\leq u_{i}(p)\,\}$
and $p=1-\alpha$ (we use $p=0.95$). Lower ACE indicates better calibration, $p>0$ indicates the prediction is under-confident, and $p<0$ indicates it is over-confident. Note that one disadvantage of ACE is that it does not penalize excessively wide intervals (e.g., under-confident intervals that always cover $s_{i}$ incur a maximum $0.05$ penalty when $p=0.95$) as much as narrow ones. Hence, we treat ACE as a secondary, human-readable diagnostic and rely on CRPS for model selection.
6 Results
In this section, we address three central questions: 1) Is instance-level performance prediction of continuous, reference-based metrics for long-form LLM outputs feasible? 2) How much supervision is necessary to achieve reasonable performance? and 3) How do confidence-based regression and reference-free LLM-as-a-Judge (Section 4) compare? We evaluate all baselines across 11 datasets and multiple metrics (Section 5.1) using RMSE to judge point estimate accuracy and CRPS to judge calibration of predictive distributions, with ACE and PC used for additional diagnostics (Section 5.2). Our key findings are as follows.
Performance prediction is indeed feasible: regression on off-the-shelf confidence estimates achieves an average RMSE of 0.14 and CRPS of 0.08, with intervals calibrated to within 16% of nominal coverage and requiring as few as 16 training examples to converge. Tables 1 and 2 show that regression on the consistency-based “DegMat” confidence estimate performs best in terms of both RMSE and CRPS.
Only 16 training instances are required for the CE-Reg’s test loss to stabilize. As long-form gold references can be expensive to collect, it is important to understand how sample-efficient these performance prediction methods would be. To answer that, we vary the size of the training set from as few as 4 instances and doubling up to the full training data size for CE-Reg. For RF-LLMaaJ we limit up to 16 instances. CE-Reg achieves close to our best observed accuracy with remarkably few labeled examples: Figure 2 shows that RMSE converges after just 16 training instances across all datasets. Even when matched to RF-LLMaaJ’s maximum of 16 in-context examples (Table 4), CE-Reg delivers substantially lower error. Although performance varies slightly across datasets, the convergence point is consistent, demonstrating that this regression approach to performance prediction can be effectively trained with a modest sample size, making it easy to collect training data even for out-of-domain tasks and mitigating potential domain differences.
CE-Reg outperforms RF-LLMaaJ. In Tables 1 and 2, CE-Reg methods achieve RMSEs averaging in the 0.14–0.15 range and reaching 0.35 at worst, depending on the task metric. Meanwhile, RF-LLMaaJ incurs much larger errors (averaging around 0.29, 0.85 at worst). Similarly, CE-Reg attains CRPS values between 0.07–0.12 averaged over task metrics, compared to about 0.21 for RF-LLMaaJ. Interestingly, most methods exhibit an over-confidence (intervals are too narrow) in ACE (95%) (Appendix 8), with verbalized CE-Reg producing ACEs of –0.89 and –0.79 and RF-LLMaaJ yielding ACEs of –0.51 and –0.52, while only some of the consistency-based features come close to nominal coverage (e.g. DegMat ACE: –0.16). Notably, this pattern largely reflects our use of RandomForest regression: substituting a XGBoost model markedly improves ACE, albeit at the expense of other performance metrics (see Figure 3 in Appendix). In addition, Judge scores and human ratings in the BiGGen-Bench-Results dataset yield similar performance, indicating our baselines generalize beyond evaluation metrics.
Prediction error is primarily metric-dependent. Across most tasks and methods, CodeBLEU and BERTScore are the easiest metrics to predict: CE-Reg consistently achieves low RMSE and CRPS and high PC on these metrics, and RF-LLMaaJ exhibits a similar trend on most datasets. The G-Eval in text summarization tasks are consistently low in RMSE and CRPS but also low in PC – this can be explained by the fact that its target score distribution is tightly concentrated. CodeJudge and LLMEvals tend to be the most difficult with high RMSE and CRPS and relatively low PC. Variation in RMSE and CRPS across evaluation metrics shows that prediction difficulty is largely metric dependent, with low metric-to-metric variability within each task. In our findings, we do not observe a similar trend for task-to-task variability. Therefore, we do not conclude anything about relative inherent difficulty between tasks.
Graph-based consistency features consistently outperform all other CE-Reg methods across every dataset and metric Among all CE-Reg, the four graph-based consistency measures (DegMat, Eccentricity, EigVal, LexSim) achieve the lowest CRPS, indicating that capturing variability across multiple generations provides a stronger predictive signal than one- or two-step verbalized confidence prompts, on all types of dataset or metrics. By contrast, the NumSemSets consistency metric performs on par with the verbalized methods, trailing the graph-based techniques. Better results from graph-based consistency over NumSemSets and verbalized confidence are more obviously reflected in the ACE and PC measures (Tables 8 and 9).
RF-LLMaaJ is a poor performance predictor. Both point prediction and interval estimation are ill-calibrated compared to CE-Reg methods. In fact, only on a few tasks does RF-LLMaaJ seem to have any ability to predict evaluation metrics at all – Appendix Table 9 shows near-zero correlation between predicted and actual scores for the majority of task metrics. Furthermore, increasing in-context examples from 4 to 16 has negligible benefit. There are some task metrics where RF-LLMaaJ significantly better than random chance, but the variability across tasks and metrics highlights how sensitive and unreliable this approach can be, relative to the consistent performance seen in the consistency-based confidence regression approach. One reason that RF-LLMaaJ struggles to predict evaluation metrics could be its sensitivity to the prompt. A key insight is that, despite both using in-prompt LLM predictions, verbalized confidence is overall more reliable than RF-LLMaaJ, which is prompted to predict arbitrary task metrics. This may reflect the LLM’s pre-training data, which likely contains the general notion of confidence much more so than task-specific metrics like CodeBLEU or GEMBA; and the few-shot examples are not enough to overcome this deficiency.
CE-Reg: Verbalized RF-LLMaaJ
2S 4-shot 8-shot 16-shot
RMSE 0.26 0.32 0.32 0.34
CRPS 0.21 0.22 0.22 0.24
Table 3: Direct Assessment: For BiGGen we evaluate two baselines using pre-generated outputs—Verbalized 2S and RF-LLMaaJ—for directly predicting human ratings. Consistency-based (requiring additional outputs) and the Verbalized 1S (requiring prompt re-design) are omitted.
CE-Reg: DegMat RF-LLMaaJ
4 inst. 8 inst. 16 inst. 4-shot 8-shot 16-shot
RMSE 0.17 0.18 0.17 0.23 0.22 0.22
CRPS 0.10 0.10 0.10 0.14 0.14 0.14
Table 4: Comparison of RMSE and CRPS for CE-Reg using DegMat feature versus reference-free LLM-as-a-Judge (RF-LLMaaJ) using different number of training examples.
(a) RMSE vs. Training Data Instances
(b) CRPS vs. Training Data Instances
Figure 2: Impact of Training Data Size on CE-Reg Performance and Test Set RMSE and CRPS Across Datasets, using the selected task-specific metric (see Table 6). The plot shows RMSE and CRPS values for training data sizes capped at 64 instances. Notably, both stabilize after 16 instances, suggesting a performance plateau for certain datasets
7 Conclusion
This work introduces the task of instance-level performance prediction for long-form LLM generation, in which a black-box model’s input–output pair is used to predict arbitrary continuous, reference-based metric scores with associated prediction intervals. To support progress, we release a comprehensive benchmark spanning 11 datasets across summarization, machine translation, code generation, and question-answering, with multiple task evaluation metrics to predict.
Our experiments compare two families of performance prediction approaches as baselines. The first–confidence-based regression–extracts off-the-shelf verbalized and consistency-based confidence signals, then trains classic regression models to output both point estimates and prediction intervals. The second–a reference-free LLM-as-a-Judge–asks an LLM, via in-context prompting, to yield metric estimates and 95% intervals for each input-output-instance. We find that confidence-based regression, particularly when using consistency features, yields substantially better RMSE and CRPS than our implemented LLM-based judging. This demonstrates that lightweight supervised regression of simple confidence estimates can outperform few-shot LLM judging for both accuracy and uncertainty quantification.
Our work lays the groundwork for research in uncertainty-aware, per-instance continuous-scored performance prediction. We hope our benchmark, code, and baselines will encourage community efforts toward robust, uncertainty-aware evaluation and deployment of long-form LLMs.
Acknowledgments
This research was supported in part by Cisco, Amazon, and by Good Systems [^1] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies). We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Appendix
A.1 Reproducibility Details
Training Regression Model Training regression models requires a set of annotated data for supervision. For a given task, a practitioner should produce a training set by determining a set of $N$ questions and gold standard responses to those questions. Then an LLM of the practitioner’s choosing should generate responses to those questions, which are compared against the gold standard responses using a task-specific evaluation metric (such as BERTScore (Zhang* et al., 2020)) – resulting in a task-specific score $s$ for each question. Additionally, the practitioner should use whatever confidence estimation methods (at their discretion) to produce confidence estimates $C$ for each question. The next step is to train a regression model of choice on this training set of $\mathcal{D}$ and $C$ over $N$ questions. This model can subsequently be used to yield a predictive distribution around the task-specific metric score for any new LLM responses to out-of-sample questions.
A.2 Sample Prompts
Prompts for Confidence (Verbalized 1S) (Tian et al., 2023)
[…]
Prompts for Confidence (Verbalized 2S) (Tian et al., 2023)
[…]
Prompts for RFLLMaaJ Baseline
[…]
Prompts for LLM-based Evaluation Metrics (G-Eval) (Liu et al., 2023)
      You will be given a news article. You will then be given one summary written for this article.        Your task is to rate the summary on one metric.        Please make sure you read and understand these instructions carefully.    Please keep this document open while reviewing, and refer to it as needed.        Evaluation Criteria:        {evaluation_instructions}        Evaluation Steps:        1. Read the news article carefully and identify the main facts and key points.    2. Read the summary and compare it to the article. Check if the summary meets the {metric} criterion.    3. Assign a score for {metric} based on the Evaluation Criteria.        Source Text:        {source}        Summary:        {prediction}            Evaluation Form (scores ONLY), provide your response in the following format:    - {metric}:        evaluation_criteria = {        "Coherence": """Coherence (1-5) - the collective quality of all sentences. A well-structured summary should build from sentence to a coherent body of information about a topic.""",        "Consistency": """Consistency (1-5) - the factual alignment between the summary and the        source document. A factually consistent summary contains only statements entailed by the source document and avoids hallucinated facts.""",        "Fluency": """Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.            - 1: Poor. The summary has many errors that make it hard to understand or sound unnatural.        - 2: Fair. The summary has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible.        - 3: Good. The summary has few or no errors and is easy to read and follow.""",        "Relevance": """Relevance (1-5) - selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries which contained redundancies and excess information."""    }
Prompts for LLM-based Evaluation Metrics (GEMBA_ref) (Kocmi and Federmann, 2023)
[…]
Prompts for LLM-based Evaluation Metrics (LLM-Eval for QA) (Xian et al., 2025; Lin and Chen, 2023)
      Task: Score the following LLM output of a question-answering task with respect to the following aspects using a 1 to 5 star rating system.            Dataset: The dataset is a Question-Answering dataset, specifically designed for evaluating factual precision and detailed comparative reasoning in AI-generated answers.            Output: Begin your evaluation by providing a short explanation. Be as objective as possible. After your explanation, provide your scores in JSON format like: [[SCORE]] {{"accuracy": 2, "informativeness": 3}}            Criteria:        - Accuracy:            1 star: Incorrect information            2 stars: Partially correct information            3 stars: Half correct information            4 stars: Mostly correct information            5 stars: Perfectly correct information            - Informativeness:            1 star: No or irrelevant information            2 stars: Very little information            3 stars: Some information            4 stars: Enough information            5 stars: Highly informative            Question:        {source}            Provided Answer:        {ans}            Reference Answer(s):        {ref_text}            Evaluation:        [[SCORE]] {{\"accuracy\": <1–5>, \"informativeness\": <1–5>}}
Prompts for LLM-based Evaluation Metrics (LLM-as-a-Judge for BiGGen) (Kim et al., 2025)
System Prompt:
      You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
User Prompt:
      ###Task Description:    An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing an evaluation criteria are given.    1. Write a detailed feedback that assesses the quality of the response strictly based on the given score rubric, not evaluating in general.    2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric.    3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"    4. Please do not generate any other opening, closing, and explanations.        ###The instruction to evaluate:    {instruction}        ###Response to evaluate:    {response}        ###Score Rubrics:    {rubric}        ###Feedback:
Prompts for LLM-based Evaluation Metrics (CodeJudge for code generation and function-calling) Tong and Zhang (2024)
Functional Correctness:
[…]
Inconsistency Prompt:
      You will be provided with a problem statement, a code snippet that supposedly addresses the problem,    and a catalog of code inconsistencies.    Evaluation Steps:    1. Read the problem statement carefully to identify the functionalities required for the    implementation.    2. Read the code snippet and compare it to the problem statement. Check if the code snippet covers    the required functionalities.    3. Output your answer in a JSON format list.    a) If the code snippet is correct, output: ["inconsistency": "None", "severity": "Negligible"].    b) If the code snippet is incorrect, output the identified inconsistencies and their severity    according to the catalog of code inconsistencies. For example: ["inconsistency": "<inconsistency1>",    "severity": "<severity1>", "inconsistency": "<inconsistency2>", "severity": "<severity2>", ...]    Problem: {PROBLEM}    Code Snippet: {CODE}    Taxonomy of Common Inconsistencies:    1. Missing dependency declarations: Negligible    2. No error messages for unexpected input cases: Negligible    3. Inefficiency, unnecessary statements: Negligible    4. Edge case not handled: Small    5. Logic error: Major    6. Function or variable not defined: Fatal    7. Code not completed: Fatal    Evaluation Form:    JSON output (a JSON list only):
A.3 Additional Dataset Information
Dataset Type Data Split Size Task-Specific Metrics
APIGen (Liu et al., 2025) Function Calling 8955 / 2985 / 2985 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
ASQA (Stelmakh et al., 2022) Question Answering 1689 / 563 / 564 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite], BERTScore
BigCodeBench (Zhuo et al., 2025) Code Generation 1994 / 665 / 665 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
BiGGen (Kim et al., 2025) Open-Ended Generation 847 / 283 / 283 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
BiGGen-Bench-Results (Kim et al., 2025) Open-Ended Generation 1662 / 555 / 555 Human Rating
Eli5 (Fan et al., 2019) Question Answering 2657 / 886 / 886 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
FActScore-Bio (Min et al., 2023) Fact-Checking 1216 / 406 / 406 FActScore [gpt-3.5-turbo]
MedLFQA (Hosseini et al., ) Medical QA 8855 / 2952 / 2952 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
Opus-100 (EN→ES) (Zhang et al., 2020; Tiedemann, 2012) Machine Translation 3579 / 1194 / 1194 BERTScore, GEMBA_SQM_ref [gemini-2.0-flash-lite]
Summeval (Fabbri et al., 2021) Text Summarization 2809 / 937 / 937 BERTScore, G-Eval (Coherence, Consistency, Fluency, Relevance) [gemini-2.0-flash-lite]
TruthfulQA (Lin et al., 2022) Question Answering 1455 / 486 / 486 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
Table 5: Our study and benchmark cover twelve datasets spanning a range of diverse tasks: code generation, fact-checking, function-calling, machine translation (MT), question-answering (QA), summarization, and others. The original dataset sizes is capped at 5,000 instances (through random sampling), with each instances generating three outputs, we collect maximum 15,000 output-score instance per datasets. Train-dev-test is split by 3:1:1. For each dataset, we also specify one or more task-specific, multi-dimensional and partial-credit evaluation metric used to evaluate generation performance.
Dataset Selected Metrics
APIGEN (Liu et al., 2025) CodeBLEU
ASQA (Stelmakh et al., 2022) BERTScore
BigCodeBench (Zhuo et al., 2025) CodeBLEU
BiGGen (Kim et al., 2025) Judge
Eli5 (Fan et al., 2019) BERTScore
FActscore-Bio (Min et al., 2023) FActScore
MED LF QA (Hosseini et al., ) BERTScore
Opus-100 (EN to ES) (Zhang et al., 2020; Tiedemann, 2012) BERTScore
Summeval (Fabbri et al., 2021) BERTScore
TruthfulQA (Lin et al., 2022) BERTScore
Table 6: While Table 5 includes our full set of task-specific evaluation metrics, for brevity we sometimes select only a single metric per dataset to report, as listed in this table here. These selected metrics are the ones reported on in Figure 2.
A.4 Complementary Tables and Figures
Dataset Task-Specific Metric Score
APIGen CodeBLEU 0.32
Judge(Func.) 0.77
Judge(Incon.) 0.56
ASQA BERTScore 0.89
LLMEval(A.) 0.87
LLMEval(I.) 0.84
BiGGen Judge 0.67
BigCodeBench CodeBLEU 0.11
Judge(Func.) 0.77
Judge(Incon.) 0.63
Summeval G-Eval(Coh.) 0.88
G-Eval(Con.) 0.95
G-Eval(Flu.) 0.59
G-Eval(Rel.) 0.79
BERTScore 0.90
Eli5 BERTScore 0.85
LLMEval(A.) 0.72
LLMEval(I.) 0.63
FActScore-Bio FActScore 0.43
Med-LFQA BERTScore 0.84
LLMEval(A.) 0.76
LLMEval(I.) 0.59
Opus-100 GEMBA-SQM_ref 0.87
BERTScore 0.85
TruthfulQA BERTScore 0.87
LLMEval(A.) 0.60
LLMEval(I.) 0.58
Table 7: Llama 3.2 11B performance on 10 long-form datasets with various task-specific evaluation metrics. All scores are normalized by their respective scale to [0,1]. Performance prediction then seeks to predict these gold-standard evaluation scores.
(a) RMSE for regression models
(b) CRPS for regression models
(c) ACE(95%) for regression models
(d) Pearson correlation for regression models
Figure 3: Comparison across regression models on development sets. For CRPS, RMSE, ACE(95%) lower is better; for Pearson correlation higher is better.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shots 16-shot APIGen Judge(Func.) -0.09 -0.53 -0.49 -0.36 -0.95 -0.94 -0.93 -0.36 -0.34 -0.34 Judge(Incon.) -0.14 -0.56 -0.52 -0.42 -0.89 -0.94 -0.86 -0.59 -0.60 -0.56 CodeBLEU -0.09 -0.54 -0.48 -0.36 -0.95 -0.94 -0.92 -0.37 -0.35 -0.36 ASQA LLMEval(A.) -0.18 -0.66 -0.15 -0.16 -0.95 -0.87 -0.82 -0.37 -0.39 -0.38 LLMEval(I.) -0.21 -0.68 -0.24 -0.23 -0.95 -0.89 -0.81 -0.41 -0.40 -0.46 BERTScore -0.24 -0.68 -0.22 -0.27 -0.90 -0.86 -0.79 -0.34 -0.33 -0.38 BiGGen Judge -0.15 -0.77 -0.26 -0.23 -0.89 -0.80 -0.68 -0.52 -0.51 -0.51 BigCodeBench Judge(Func.) -0.20 -0.37 -0.20 -0.18 -0.92 -0.94 -0.78 -0.39 -0.42 -0.40 Judge(Incon.) -0.14 -0.45 -0.17 -0.13 -0.95 -0.93 -0.92 -0.49 -0.52 -0.49 CodeBLEU -0.19 -0.48 -0.23 -0.22 -0.91 -0.90 -0.76 -0.43 -0.43 -0.38 Eli5 LLMEval(A.) -0.20 -0.91 -0.22 -0.21 -0.87 -0.92 -0.83 -0.36 -0.35 -0.33 LLMEval(I.) -0.18 -0.90 -0.19 -0.17 -0.95 -0.93 -0.76 -0.22 -0.22 -0.18 BERTScore -0.22 -0.86 -0.22 -0.22 -0.88 -0.85 -0.76 -0.63 -0.57 -0.55 Fact-Bio FActScore -0.14 -0.86 -0.13 -0.11 -0.90 -0.82 -0.81 -0.40 -0.39 -0.38 MED LF QA LLMEval(A.) -0.18 -0.84 -0.19 -0.19 -0.95 -0.85 -0.84 -0.37 -0.34 -0.33 LLMEval(I.) -0.17 -0.84 -0.19 -0.18 -0.59 -0.95 -0.65 -0.29 -0.30 -0.29 BERTScore -0.19 -0.83 -0.22 -0.22 -0.93 -0.90 -0.85 -0.49 -0.45 -0.45 Opus-100 GEMBA -0.17 -0.45 -0.30 -0.24 -0.94 -0.84 -0.80 -0.44 -0.40 -0.38 BERTScore -0.10 -0.45 -0.25 -0.21 -0.91 -0.89 -0.80 -0.40 -0.34 -0.32 Summeval G-Eval(Coh.) -0.18 -0.88 -0.18 -0.20 -0.95 -0.94 -0.92 -0.78 -0.81 -0.86 G-Eval(Con.) -0.06 -0.82 -0.07 -0.07 -0.95 -0.95 -0.87 -0.44 -0.40 -0.40 G-Eval(Flu.) 0.04 -0.82 0.04 0.04 -0.95 -0.92 -0.10 -0.29 -0.26 -0.24 G-Eval(Rel.) -0.17 -0.87 -0.19 -0.19 -0.95 -0.94 -0.93 -0.61 -0.58 -0.57 BERTScore -0.20 -0.86 -0.20 -0.21 -0.90 -0.89 -0.81 -0.51 -0.52 -0.50 TruthfulQA LLMEval(A.) -0.16 -0.70 -0.25 -0.20 -0.88 -0.92 -0.79 -0.38 -0.35 -0.37 LLMEval(I.) -0.23 -0.55 -0.29 -0.28 -0.63 -0.76 -0.74 -0.44 -0.43 -0.42 BERTScore -0.19 -0.80 -0.28 -0.25 -0.86 -0.81 -0.71 -0.52 -0.50 -0.50 AVERAGE -0.16 -0.70 -0.23 -0.21 -0.90 -0.89 -0.79 -0.44 -0.43 -0.42
Table 8: ACE(95%) Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings. For ACE(95 %), values closer to zero are better. Negative values indicate intervals that are too narrow (over-confidence), while positive values indicate intervals that are too wide (under-confidence). Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.
CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat. Eccen. EigVal. LexSim. NumSet. 1S 2S 4-shot 8-shots 16-shot APIGen Judge(Func.) 0.48 0.38 0.38 0.42 0.34 0.19 0.16 0.25 0.27 0.23 Judge(Incon.) 0.55 0.35 0.30 0.35 0.17 0.09 0.13 0.03 0.01 0.08 CodeBLEU 0.65 0.38 0.41 0.47 0.08 0.02 0.11 0.05 0.04 0.01 ASQA LLMEval(A.) 0.47 0.30 0.43 0.52 0.23 0.41 0.35 0.14 0.14 0.11 LLMEval(I.) 0.38 0.27 0.36 0.42 0.16 0.33 0.27 0.02 0.05 0.08 BERTScore 0.51 0.35 0.53 0.58 0.07 0.34 0.16 0.07 0.05 0.10 BiGGen Judge 0.62 0.26 0.52 0.55 0.10 0.21 0.25 0.26 0.21 0.06 BigCodeBench Judge(Func.) 0.46 0.36 0.41 0.48 0.07 0.11 0.15 0.06 0.01 0.02 Judge(Incon.) 0.33 0.27 0.26 0.35 0.04 0.08 0.13 0.09 0.02 0.04 CodeBLEU 0.56 0.47 0.50 0.56 0.09 0.00 0.05 0.07 0.03 -0.07 Eli5 LLMEval(A.) 0.50 0.18 0.48 0.46 0.31 0.16 0.21 0.12 0.09 0.15 LLMEval(I.) 0.36 0.13 0.31 0.30 0.23 0.17 0.22 0.22 0.24 0.18 BERTScore 0.54 0.32 0.53 0.51 0.06 0.01 0.11 -0.02 -0.04 0.02 Fact-Bio FActScore 0.72 0.25 0.74 0.77 0.27 0.68 0.26 0.47 0.50 0.54 MED LF QA LLMEval(A.) 0.48 0.24 0.49 0.50 0.24 0.31 0.24 0.11 0.13 0.13 LLMEval(I.) 0.44 0.30 0.42 0.43 0.07 0.28 0.24 0.24 0.28 0.27 BERTScore 0.55 0.38 0.52 0.54 0.16 0.22 0.36 0.09 0.04 0.05 Opus-100 GEMBA 0.52 0.37 0.48 0.55 0.29 0.24 0.14 0.09 0.11 0.08 BERTScore 0.66 0.50 0.63 0.65 0.23 0.11 0.15 0.06 0.05 -0.02 Summeval G-Eval(Coh.) 0.38 0.17 0.36 0.35 0.09 0.23 0.02 0.00 0.04 0.02 G-Eval(Con.) 0.29 0.10 0.34 0.30 0.12 0.17 0.02 -0.02 -0.04 -0.07 G-Eval(Flu.) 0.08 0.00 0.08 0.08 0.08 0.04 -0.03 -0.04 0.05 -0.03 G-Eval(Rel.) 0.40 0.13 0.38 0.38 0.16 0.23 0.02 0.02 0.03 0.01 BERTScore 0.56 0.24 0.57 0.52 0.08 0.16 0.00 -0.03 0.04 -0.02 TruthfulQA LLMEval(A.) 0.60 0.24 0.54 0.54 0.12 0.18 0.31 0.06 -0.01 0.12 LLMEval(I.) 0.43 0.21 0.35 0.39 0.14 0.20 0.18 0.15 0.06 0.20 BERTScore 0.60 0.13 0.39 0.46 0.08 0.12 0.11 0.03 0.10 -0.02 AVERAGE 0.49 0.27 0.43 0.46 0.15 0.20 0.16 0.10 0.09 0.08
Table 9: PC Benchmark. We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings. For PC, values higher the better. Judge(Func.) = CodeJudge Functional Correctness. Judge(Incon.) = CodeJudge Inconsistency. LLMEval(A.) = LLM Eval Accuracy. LLMEval(I.) = LLMEval Informativeness.

### v2

@@
-Instance-level Performance Prediction for Long-form Generation Tasks
+PIE: Performance Interval Estimation for Free-Form Generation Tasks
 Abstract
-We motivate and share a new benchmark [^1] Benchmark URL forthcoming for instance-level performance prediction of long-form generation tasks having multi-faceted, fine-grained quality metrics.
-Our task-, model- and metric-agnostic formulation predicts continuous evaluation metric scores given only black-box model inputs and outputs.
-Beyond predicting point estimates of metric scores, the benchmark also requires inferring prediction intervals to quantify uncertainty around point estimates.
-Evaluation spans 11 long-form datasets/tasks with multiple LLMs, baselines, and metrics per task.
-We show that scores can be effectively predicted across long-form generation tasks using as few as 16 training examples Overall, we introduce a novel and useful task, a valuable benchmark to drive progress, and baselines ready for practical adoption today.
+Confidence estimation infers a probability for whether each model output is correct or not.
+While predicting such binary correctness is sensible for tasks with exact answers, free-form generation tasks are often more nuanced, with output quality being both fine-grained and multi-faceted.
+We thus propose Performance Interval Estimation (PIE) to predict both: 1) point estimates for any arbitrary set of continuous-valued evaluation metrics; and 2) calibrated uncertainty intervals around these point estimates.
+We then compare two approaches: LLM-as-judge vs.
+classic regression with confidence estimation features.
+Evaluation over 11 datasets spans summarization, translation, code generation, function-calling, and question answering.
+Regression is seen to achieve both: i) lower error point estimates of metric scores; and ii) well-calibrated uncertainty intervals.
+To support reproduction and follow-on work, we share our data and code.
+[^1] Data and code will be shared upon paper acceptance.
 **footnotetext: Both authors contributed equally.****footnotetext: Work does not relate to the author’s position at Amazon.
 1 Introduction
-Large language models (LLMs) can tackle increasingly challenging open-ended, long-form generation tasks across diverse modalities (e.g., text, images, code, etc.) Liu et al.
-(2024); Yin et al.
+Large language models (LLMs) are fallible: the quality of their outputs varies substantially across inputs, tasks, models, and evaluation metrics Schellaert et al.
 (2025).
-Given the vast output space, many different generations of varying quality are possible Giulianelli et al.
-(2023).
-This motivates evaluation of long-form output quality that is: 1) fine-grained and nuanced (rather than binary correct/incorrect), and 2) multi-faceted, with diverse evaluation metrics used to assess different aspects of output quality Gao et al.
-(2025).
-For example, summarization quality might be measured via aspects such as brevity, fluency, coherence, consistency, and relevancy.
-Like other AI models, LLMs are fallible, with varying output quality across different inputs Schellaert et al.
-(2025).
-Consequently, the ability to predict the quality of each output – across different inputs, tasks, models, and metrics – would have myriad benefits.
-Key among these is to guide appropriate trust and use for both human users Steyvers et al.
+As LLMs are increasingly applied to real-world tasks, the ability to predict the quality of each individual output is important to guide appropriate trust in model outputs for both human users and downstream agents Steyvers et al.
 (2025); Dhuliawala et al.
 (2023); Bhatt et al.
-(2021) and downstream agents Mielke et al.
+(2021); Mielke et al.
 (2022); Yoffe et al.
 (2024).
-For example, outputs predicted to be of high quality might be used as-is, while those of lesser quality might be revised or discarded accordingly.
-Predicted quality could also provide a signal to optimize prompt selection or model parameterization Zhou et al.
-(2022), to select between different generations Chen et al.
-(2023), or to route inputs to whichever model is predicted to perform best Lee et al.
+Confidence estimation (CE) canonically estimates the probability that a model output is correct or not.
+Predicting such binary correctness is sensible for tasks with exact answers, but modern free-form generation tasks are often more nuanced (e.g., summarization, translation, code generation, etc.).
+With such tasks, output quality often varies both 1) continuously; and 2) along multiple aspects (aka dimensions or properties), depending on the task.
+Example aspects might include properties such as fluency, relevancy (Liu et al., 2023), factual consistency (Min et al., 2023), etc.
+Predicting only unidimensional binary correctness can thus fail to capture more nuanced differences and other important dimensions of output quality.
+Why represent output quality as continuous?
+First and foremost, fine-grained evaluation metrics typically yield continuous scores; we want to be able to predict any arbitrary metric.
+Second, while subjective human quality ratings may be labeled ordinally, more objective quality measures are naturally continuous (e.g., execution time of generated code or time required to further revise model outputs).
+Framing the quality prediction task as regression allows us to be agnostic as to both: 1) whether the prediction target is ordinal or continuous; and 2) whether we are predicting an observed value (e.g., human rating or program execution time) or a derived evaluation metric.
+This regression framing also has important implications for uncertainty quantification (UQ).
+When quality is binary, we can formulate uncertainty as the probability of an output being correct or not.
+However, when output quality is continuous, UQ must also shift to estimating confidence intervals around point estimates of output quality.
+Such UQ is particularly consequential in high-stakes settings where decisions are made directly from model outputs, necessitating risk-aware judgments.
+For example, consider software development workflows where LLM-generated code is increasingly executed or deployed with limited human oversight (Cotroneo et al., 2025; Navneet and Chandra, 2025).
+High confidence is required that generated code will reliably execute, pass unit tests, and be efficient and robust (Tong and Zhang, 2024) [^2] Predicting generated code quality pre-execution avoids both potential security risks and any time/space requirements..
+To enable risk-informed decision-making, a point estimate of output quality alone is insufficient; a confidence interval helps lower-bound quality to better assess the level of risk.
+For example, imagine generated code has predicted functional correctness of 0.89 with a narrow 95% interval (e.g., $\pm 0.09$), meaning 95% confidence that output quality is $\geq$ 80%.
+Assuming point estimates and confidence intervals are well-calibrated (i.e., trustworthy), this may be considered sufficiently low risk to use the output without human review.
+Alternatively, if the same predicted correctness of 0.89 were instead accompanied by a wide interval (e.g., $\pm 0.28$), this would indicate substantial uncertainty and a more serious worst-case scenario: i.e., 95% confidence that output quality is only $\geq$ 0.61.
+In practice, such uncertainty can translate into failures that are costly to detect and resolve, motivating additional safeguards, such as requiring human review and revision.
+Table 1 illustrates such an example in more detail.
+Prediction intervals are thus essential for guiding risk-aware decision making (Hullman et al., 2025).
+Intervals must also be well-calibrated, since miscalibrated intervals can do harm in leading users astray.
+In sum, this work identifies the need for: 1) supporting arbitrary, task-specific measures of output quality for nuanced, free-form generation tasks; 2) continuous, multidimensional point estimates of output quality; and 3) confidence intervals around point estimates.
+To address these needs, we propose Performance Interval Estimation (PIE).
+The goal of PIE is to enable calibrated trust in model outputs for such tasks via a combination of predicting: 1) point estimates of continuous, task-specific quality measures; and 2) calibrated confidence intervals around each point estimate.
+By jointly modeling quality and uncertainty, PIE supports practical deployment decisions such as abstention, asking clarification questions, or fallback to alternative models in low-confidence cases (Aliannejadi et al., 2019; Wen et al., 2025; Chuang et al., 2025).
+High-quality outputs may be accepted directly, while low-quality ones can be revised, regenerated, or discarded.
+Predicted quality could further support prompt or model selection Zhou et al.
+(2022), candidate ranking Chen et al.
+(2023), and routing inputs to models expected to perform best Lee et al.
 (2024); Šakota et al.
@@
 (2023); Ding et al.
-.
-Assuming evaluation metrics for fine-grained output quality are continuous $\in[0,1]$, we not only want to predict point estimates for metric scores, but also prediction intervals to quantify uncertainty around point estimates Shrestha and Solomatine (2006); Patel (1989); Geisser (1993).
-For example, given a point estimate prediction of 0.8, a tight 95% interval (say [0.79, 0.81]) would predict that output quality is quite close to the point estimate, while a loose 95% interval (say [0.6, 0.1]) would mean that quality may actually be much lower.
-Prediction intervals thus provide crucial uncertainty quantification for risk-averse settings by indicating probabilistic lower bounds on expected performance Harrington et al.
 (2024).
-We thus emphasize the importance of prediction intervals as well as point estimates for performance prediction.
-In this work, we motivate and formulate the task of instance-level performance prediction for long-form generation tasks: predicting instance-level output quality – point estimates and prediction intervals – for continuous evaluation metrics.
-By design, our task formulation is task-, model-, and metric-agnostic: given an input-output pair for any black-box model, one must predict arbitrary evaluation metrics.
-The generality of this formulation supports broad use across the diverse, ever-changing landscape of LLM models and evaluation metrics for long-form output quality.
-We assume supervision is key: by observing examples of past input-output pairs and corresponding metric scores, one can learn a metric-specific prediction function.
-An ideal solution would achieve strong predictive performance with minimal supervision.
-To promote research on instance-level performance prediction for long-form generation tasks, we create and share a new benchmark spanning twelve long-form tasks/datasets: text summarization, machine translation, code generation, and question answering.
-With different metrics often used to assess different aspects of output quality, our benchmark includes multiple evaluation metrics for most tasks.
-In regard to LLM coverage , we include outputs from Llama 3.2 ;other LLMs could be easily added later.
-Our benchmark includes two baselines.
-One uses off-the-shelf confidence estimators (verbalized confidence and output consistency) to elicit weak features, then trains a classic regression model (e.g., linear regression, random forests) to predict the metric score from the features.
-The other baseline uses a reference-free LLM-as-a-Judge Li et al.
-(2024a) with in-context learning for textual regression Vacareanu et al.
-(2024) to directly predict the metric score from the input-output pair.
-Our work demonstrates the feasibility of instance-level prediction of continuous evaluation metric scores and associated uncertainty intervals.
-We show that scores can be predicted with low RMSE and CRPS, across long-form generation tasks, and using as few as 16 training examples (Section 5.2).
-While the feature-based regression baseline described above is shown to perform well, the LLM judge baseline proves inconsistent.
-We also see that some confidence-based features are much more predictive than others.
-Taken together, our formulation, benchmark, and findings provide both a strong foundation for future research as well as baselines suitable for practical use today.
+For reproducible benchmarking on PIE’s prediction objective, we identify a set of 11 datasets spanning summarization, machine translation, function-calling, code generation, and question answering.
+Since different evaluation metrics capture different aspects of output quality, we include multiple metrics per task.
+To benchmark over outputs from different generative models, we include outputs from Llama 3.2-11B and Gemini 2.0-flash-lite.
+We then evaluate two approaches: reference-free LLM-as-judge vs.
+classic regression using confidence estimation features.
+For the former, we prompt an LLM to directly predict the target evaluation metric score and a corresponding prediction interval for each generated output.
+Our evaluation includes use of in-context learning with a small set of labeled examples.
+The classic regression model predicts metric scores and intervals from confidence estimation signals, including verbalized confidence, consistency-based measures, and other uncertainty methods, by training a regression model to map these features to continuous metric targets.
+Results across the 11 datasets, 15 metrics, and 2 LLMs show that the regression-based approach performs best in estimating both point estimates and confidence intervals for metric scores.
+Because free-form gold references can be expensive to collect, sample-efficiency is also important.
+Results show the regression-based approach converges after only 16 training instances across all 11 datasets.
+Overall, we conceptually integrate three theoretically related but methodologically distinct lines of prior work: confidence estimation, quality estimation, and LLM-as-a-judge (Section 2).
+In reviewing each line’s core assumptions, methodological choices, and limitations, we show none provides a task-agnostic, general solution for predicting uncertainty in free-form generation.
+We then introduce a unified task formulation, PIE, to predict continuous evaluation metrics together with calibrated uncertainty intervals (Section 3).
+Since PIE is agnostic regarding evaluation based on directly observed values or derived performance metrics, it is equally suited to predicting both reference-based quality metrics (e.g., based on human ratings) as well as more objective metrics like code run-time, memory use, or tool call alignment.
+To enable reproducible benchmarking, we collate and share a multi-task, multi-metric benchmark dataset (Section 4).
+Experiments (Section 5) and findings (Section 6) shows that classic regression matches or exceeds contemporary LLM-as-judge methods, consistently achieving more accurate and better-calibrated intervals while also using as few as 16 labeled examples to do so (Section 6).
+Input Scale the input field to the range [0, 1] and display it as a DataFrame.
+LLM output scaler = MinMaxScaler( feature_range=(0, 1))
+scaled_values = scaler.fit_transform([[x] for x in l])
+df = pd.DataFrame(scaled_values, columns=[’Scaled Values’])
+return df
+Target text scaler = MinMaxScaler()
+l_scaled = scaler.fit_transform( l.reshape(-1, 1))
+df = pd.DataFrame(l_scaled, columns=[’Scaled Values’])
+return df
+Prediction $0.8918\pm 0.2824$ at 95% probability
+Gold score $0.7500$ (functional correctness)
+Table 1: Illustrative PIE example on BigCodeBench.
+The CE-Reg predictor estimates functional correctness and a 95% prediction interval.
+This interval covers the gold score despite an optimistic point estimate, mitigating overconfident assessment of code that could fail for array-shaped inputs.
+Table 2: Comparison of output quality estimation paradigms, illustrating their respective strengths and limitations in supporting continuous, multi-dimensional evaluation and uncertainty for free-form generation.
+Confidence Estimation Quality Estimation Reference-free LLM-as-a-Judge Performance Interval Estimation
+Used mostly in Limited (QA) Limited (MT) Various tasks Various tasks
+Continuous targets No Yes Yes Yes
+Multiple dimensions No Limited (error types) Yes Yes
+Explicit uncertainty Yes Limited Limited Yes
 2 Related Work
-Our formulation of performance prediction draws upon a large body of prior work across related areas: Confidence Estimation and Uncertainty Quantification (Section 2.1), Score and Benchmark Predictability (Section 2.2), Query Performance Prediction in information retrieval (Section 2.3), Quality & Metric Estimation in machine translation (Section 2.4), Textual Regression (Section  2.5), and LLM-as-a-judge & Reward Models (Section 2.6).
-2.1 Uncertainty and Confidence Estimation
-Our work is particularly motivated by confidence estimation (CE) Xiong et al.
-(2024) and uncertainty quantification (UQ) Vashurin et al.
-(2025); Xiao et al.
-(2022), which similarly predict how much each system output can be trusted.
-While terminology and definitions vary across different works, UQ (Smith, 2013) canonically seeks to quantify certainty regarding a given process, observation, or system output, while CE seeks to predict the probability that a system output is correct or not (Guo et al., 2017; Steyvers et al., 2025).
-However, as Huang et al.
-(2024) write, “the correctness of long-form generation is not always either true or false but can be partially correct.
-Therefore, a single confidence score for long-form outputs is ambiguous: it can either imply ‘I am x% confident that the answer is 100% correct’ or ’I am 100% confident that the answer is x% correct.” Moreover, Ulmer et al.
-(2024) suggest estimation of LLM confidence intervals as future work.
-Nevertheless, prior work has typically modeled long-form generation confidence via binary correctness by: i) mapping generations to a most similar multiple choice answer (Tjuatja et al., 2024), ii) binarizing partial-credit scores via thresholding (Kuhn et al., 2023; Lin et al., 2024; Band et al., 2024; Pedapati et al., 2024), iii) testing whether or not the output contains an answer (string) (Rajpurkar et al., 2016; Yu et al., 2018), iv) focusing on tasks with short answers wherein evaluation can rely on exact matching (Pedapati et al., 2024; Xiao et al., 2022); or v) decomposing a long output into smaller nuggets checked individually for correctness (Zhang et al., 2025, 2024a; Min et al., 2023).
-In recent work, Zhang et al.
-(2025) usefully infer partial-credit scores for correctness at various granularities (e.g.
-sentence or passage level) but still assume that output quality can be measured via a single dimension of correctness.
-Huang et al.
-(2024) also emphasize “…the multifaceted nature of evaluating long-form outputs, where factors like factuality, coherence, clarity, and comprehensiveness each play a role…” However, they assume evaluation dimensions are implicit and latent, leading to subjective disagreements between evaluators.
-In contrast, we assume distinct evaluation metrics assess different aspects of output quality, and predict each separately.
-Prior studies also tend to show only correlation between CE vs.
-evaluation metric scores, while we predict metric scores directly.
-Our formulation pursues uncertainty quantification via prediction intervals that quantify uncertainty around point estimates of metric scores.
-2.2 Score & Benchmark Predictability
-Like CE and UQ above, score predictability work Hernández-Orallo et al.
-(2022); Pacchiardi et al.
-(2025); Schellaert et al.
-(2025) similarly makes instance-level predictions of binary output correctness.
-Hernández-Orallo et al.
-note continuous prediction is more complex and proceed with binary.
-Schellaert et al.
-suggest future non-binary work while Pacchiardi et al.
-assume the binary case.
-Three other differences are noteworthy.
-First, score predictability emphasizes anticipatory prediction of instance-level success before output generation.
-This avoids computing outputs but makes prediction more difficult (without need for our use cases).
-Our performance prediction task more closely resembles UQ and CE in which confidence is inferred for a given output.
-Second, model details are often exploited to improve prediction, vs.
-our model-agnostic, black-box formulation.
-Third, score predictability methods tend to leverage a large feature set (e.g.
-text embeddings); we require only (input, output, score) triples.
-Other work on benchmark prediction infers continuous scores, but only for entire datasets in aggregate Ye et al. (2023); Zhang et al. (2024b). Instead, we predict the quality of individual outputs.
-2.3 QPP in Information Retrieval
-Our nomenclature of “performance prediction” is inspired by query performance prediction (QPP) work in information retrieval (Cronen-Townsend et al., 2002; He and Ounis, 2006; Arabzadeh et al., 2024).
-Given an input query, QPP seeks to predict the quality of output search results.
-Like our work, QPP is both instance-based and predicts arbitrary, continuous-valued metrics of output quality.
-QPP work is split into pre-retrieval vs.
-post-retrieval methods.
-Like anticipatory score predictability (Section 2.2), pre-retrieval methods require predicting output quality (of search results) before executing the (input) query.
-In contrast, post-retrieval methods run the query and analyze output search results to better predict their quality, and this helps them to perform much better in practice.
-We also expect output analysis to help prediction.
-2.4 Quality and Metric Estimation in MT
-Quality estimation (QE) in machine translation (MT) predicts translation output quality at run-time Specia et al.
+In this section, we review and compare key related work in confidence estimation (Section 2.1), quality estimation (Section 2.2), and (reference-free) LLM-as-judge (Section 2.3).
+Each of these varies in how uncertainty and metric prediction are considered.
+To relate these approaches and their relative capabilities, we compare them via several criteria in particular regard to evaluating the quality of free-form generations, as summarized in Table 2.
+2.1 Confidence Estimation (CE)
+Confidence estimation canonically models the reliability of a system output as the probability of an output being correct or not, producing a single scalar confidence value tied to binary correctness (Blatz et al., 2004; Guo et al., 2017; Steyvers et al., 2025).
+This view naturally fits short, discrete tasks with exact answers, such as factoid questions where correctness can be treated as all-or-nothing.
+Two CE approaches are particularly well-known for their simplicity and applicability to black-box models: verbalized CE Xiong et al.
+(2024); Tian et al.
+(2023) and consistency-based CE (Kuhn et al., 2023; Lin et al., 2024).
+The former prompts an LLM to self-report its confidence alongside the generated output in a single LLM invocation.
+Consistency-based methods instead infer confidence from the degree of agreement across multiple independent generations.
+Computational overhead thus scales linearly with the number of samples.
+As discussed earlier, framing output quality wrt.
+binary correctness is less suited to free-form, nuanced generation tasks in which assessment of output quality is often fine-grained (i.e., continuous), multi-dimensional, and task-dependent.
+2.2 Quality Estimation (QE)
+Quality estimation in machine translation (MT) predicts output translation quality at runtime Specia et al.
 (2009, 2010).
-Earlier work on CE for MT also assumes binary correctness Blatz et al.
-(2004) which QE then generalizes to continuous, partial-credit evaluation.
-The term CE is also still sometimes used to refer to the broader QE task Chelba et al.
+In comparison to PIE, it shares both important similarities and notable differences.
+For example, early work on CE for machine translation also assumed binary correctness Blatz et al.
+(2004), which QE then generalized to continuous, partial-credit evaluation Specia et al.
+(2009, 2010).
+PIE is similarly inspired but extends this to diverse tasks, arbitrary metrics, and a strong emphasis on uncertainty and calibrated prediction intervals.
+Work on QE has been largely restricted to MT, though other potential application areas have been discussed.
+For example, Scarton et al.
+(2016) suggested generalizing QE to tasks such as correcting grammar Napoles et al.
+(2016), generation Dušek et al.
+(2019), and others Specia et al.
 (2018).
-QE is canonically framed as a regression task to predict human quality ratings (i.e., “direct assessment”) of output quality.
-Scarton et al.
-(2016) frame QE more broadly as predicting output quality across diverse NLP tasks, including text summarization and simplification as well as QE.
-Some QE work has also emphasized the value of inferring prediction intervals beyond simple point estimates Glushkova et al.
-(2021); Giovannotti (2023); Zerva and Martins (2024).
-Zouhar et al.
-(2023) propose “metric estimation”: predicting evaluation metric scores rather than human quality ratings.
-However, their goal in doing so is to pre-train a QE model.
-We are agnostic as to the type of prediction target (i.e., metric score or human rating) used to evaluate long-form outputs.
-To illustrate this, we also include a “direct assessment” task in our evaluation (Section 5.1), complementing our other metric prediction results.
-2.5 Textual Regression
-Wang et al.
-(2022) note “almost no work in NLP on calibration in a regression setting” and consider three regression tasks: semantic similarity, sentiment analysis, and QE.
-They use BERT/SBERT and report point estimates and prediction intervals.
-LLM-based work has pursue three primary approaches: i) extracting LLM features for classic regression Tang et al.
-(2025a), ii) leveraging in-context learning Vacareanu et al.
-(2024); and iii) fine-tuning for specific regression tasks Tang et al.
-(2025b).
-Vacareanu et al.
-(2024) consider three synthetic regression tasks and show effective results across several models via in-context learning, but without prediction intervals.
-We report a similar in-context LLM baseline in this work (Section 4.3).
-Most recently, Vedula et al.
-(2025) propose quantile regression for LLM-based price estimation, reporting prediction intervals as well as point estimates.
-2.6 LLM-as-a-Judge & Reward Models
-The rise of LLMs has also brought the advent of LLM-as-a-judge.
-A key distinction exists between reference-free (RF) methods, suitable for predicting output quality at run-time, vs.
-reference-based (RB) methods that require human reference outputs to compare AI outputs against Sheng et al.
-(2024); Li et al.
-(2024a).
-Tan et al.
-(2024) present an excellent discussion relating RF methods to reward models used in reinforcement learning from human feedback (RLHF), noting “these two fields are closely related and are gradually converging.” Our work is related to both fields, though we focus on run-time prediction of RB metrics.
-We also note similarity between RF LLM judges and “verbalized” CE methods (Section 2.1) in which an LLM is prompted for its confidence that a given output is correct (Tian et al., 2023; Xiong et al., 2024).
-Both support run-time use of LLMs through prompting, though verbalized CE methods have assumed a particular type of assessment (i.e., confidence) while RF LLM judges have been conceived and pursued much more broadly.
-While RF LLM judges are increasingly being used as final evaluators, their reliability remains contested: they can exhibit self-preference and positional biases Wang et al.
-(2024); Li et al.
-(2024b), mis-rank instances they cannot themselves solve without a human reference answer Krumdick et al.
-(2025), and underperform or distort correctness in domains like software engineering Zhou et al.
+In contrast, PIE is fundamentally conceived as task- and domain-agnostic.
+Canonically, QE also refers to predicting human quality ratings, i.e., “direct assessment” judgments of translation quality.
+That said, some work has considered other metrics such as post-editing time Specia et al.
 (2018).
-More broadly, RF methods have inherent structural biases—being effectively one generation model evaluating another—which risks test-time gaming, favoring models similar to the judge, and even disadvantaging higher-quality (including human) outputs Deutsch et al.
-(2021).
-Reproducibility can also be limited by underlying changes in closed LLM judges or prompt sensitivity in use of open judges Bavaresco et al.
-(2025).
-Meta-evaluation of RF methods often reports correlation with human judgments, but caution is warranted.
-For example, the breadth and diversity of correlation studies is limited Bavaresco et al.
-(2025).
-Krumdick et al.
-(2025) note correlation can also mask disagreement on difficult instances, limiting ability to evaluate more advanced models or reliably predict instance-level quality.
-Dorner et al.
-(2025) show that when a RF LLM judge has comparable accuracy to an evaluated model, “no debiasing method can decrease the required amount of ground truth labels by more than half… [showing] severe limitations of the [RF] LLM-as-a-judge paradigm…" Their empirical results show practical performance is even worse than this theoretical limit.
-Gera et al.
-(2025) similarly report, “Crucially, even very good instance-level capabilities do not guarantee accurate model ranking.”
-While we expect RF LLM judges to both continue to improve and become better understood, the limitations above suggest that RB LLM judges (and perhaps classic RB metrics) still remain an important part of the NLP evaluation toolkit for evaluating long-form outputs.
-This motivates a continuing need for run-time prediction of RB evaluation metrics, as we pursue in this study.
-3 Task Formulation
-Prediction intervals represent a natural extension of uncertainty for point estimates, bounding the range within which the true performance score will fall with a specified probability (e.g.
-95 %) Shrestha and Solomatine (2006); Patel (1989); Geisser (1993).
-Prediction intervals are increasingly being reported, e.g., in MT quality estimation (Section 2.4) and other textual regression tasks (Section 2.5).
+In contrast, PIE is fundamentally metric-agnostic.
+While both QE and PIE are concerned with output “quality”, PIE’s notion of quality is broader, encompassing arbitrary task-dependent metrics of quality, such as the functional correctness of code.
+Another key difference between QE and PIE lies in their respective treatments of uncertainty quantification (UQ).
+While recent work has explored uncertainty within QE, principally via conformal prediction (Glushkova et al., 2021; Giovannotti, 2023; Zerva and Martins, 2024), even here, uncertainty is not treated as a core modeling objective, and the scope remains confined to translation.
+In contrast, UQ is central and fundamental to PIE to support risk-aware decision making across diverse tasks and settings.
+As Table 1 illustrates, even a high predicted quality score can still present significant risk if the prediction interval is wide.
+Within QE, Zouhar et al.
+(2023) proposed metric estimation: predicting reference-based evaluation metrics without access to the human reference.
+Similar to PIE, the focus is on predicting evaluation metrics, not just human judgments.
+However, like most QE work, its scope is translation-specific, focused on a small set of reference-based metrics, and predicts only point estimates.
+2.3 LLM-as-a-Judge (LLMaaJ)
+The rise of LLMs has also brought the LLM-as-a-judge paradigm.
+A key distinction is whether an LLMaaJ is used at runtime (online) to predict output quality vs.
+being used during evaluation (offline) to evaluate output quality.
+Closely related to this is the distinction between reference-based (RB) vs.
+reference-free (RF) methods.
+RB methods require a human (gold) reference output with which to compare the generated output and so are only suitable for use during evaluation (offline) (Sheng et al., 2024; Li et al., 2024).
+In contrast, RF methods have no such requirement, assessing output quality without comparing to any reference output (Liu et al., 2023).
+As such, RF methods are more flexible and can be used both online and offline.
+We view RF-LLMaaJ as closely related to CE and QE since all three seek to predict output quality at runtime.
+However, a key limitation of LLM approaches seen in prior studies is their limited ability to regress continuous values, often struggling to produce calibrated intervals (Vedula et al., 2025), also seen in our own findings (Section 6).
+RF-LLMaaJ shares conceptual similarities with “verbalized” CE methods (Section 4.1) that prompt a model for its confidence in its output correctness (Tian et al., 2023; Xiong et al., 2024).
+Both rely on prompting an LLM at runtime, but verbalized CE assumes a specific assessment target (confidence), whereas RF-LLMaaJ methods have been explored more broadly in evaluating model outputs.
+RF-LLMaaJ methods also share conceptual similarities with QE methods (Section 2.2).
+As discussed, QE canonically seeks to predict human ratings (for translation quality) and only rarely to predict objective measures (e.g., post-editing time Specia et al.
+(2018)) or evaluation metrics Zouhar et al.
+(2023)).
+Similarly, RF-LLMaaJ work also typically seeks to predict human judgments and is evaluated by how well it aligns with them, rather than predicting reference-based evaluation metrics.
+As noted earlier, prediction of confidence intervals is atypical in QE work (Glushkova et al., 2021; Giovannotti, 2023; Zerva and Martins, 2024).
+Similarly, uncertainty quantification for continuous predictions (e.g., inferring confidence intervals) is also rare with LLMs (Hashemi et al., 2024).
+As in CE and QE, a key premise of our work is that (offline) evaluation requires information not available (online) at runtime.
+Examples include human reference outputs or knowledge such as memory use by generated code once executed.
+The key challenge is thus online prediction of offline evaluation.
+However, when online RF-LLMaaJ is also the standard for offline evaluation, there is no need to predict it.
+We posit that despite growing popularity of evaluation via RF-LLMaaJ methods, many important evaluation settings will continue to require information not available at runtime.
+3 PIE Task Formulation
+As summarized in Table 2, existing methods either model uncertainty over binary correctness or predict continuous quality without task-general uncertainty estimates.
+We therefore propose PIE, which given a model output, predicts a continuous evaluation metric for it, along with a calibrated uncertainty interval.
+PIE extends point-estimate prediction by bounding the range within which the true performance score is expected to fall with a specified probability (e.g., 95%) Shrestha and Solomatine (2006); Patel (1989); Geisser (1993).
 Formalization.
-Let $\mathcal{X}$ denote the input space (user queries, upstream agent inputs) and $\mathcal{Y}$ the long-form output space (e.g., text, code, etc.) for both LLM generations and gold reference outputs.
-We assume each task $t\in\mathcal{T}$ has $K_{t}$ continuous, reference-based (RB) evaluation metrics $\mathcal{M}_{t}=\{m_{t,1},\dots,m_{t,K_{t}}\}$, where each metric $m_{t,k}:\mathcal{Y}\times\mathcal{G}\rightarrow\mathbb{R}$ assigns LLM output $y$ a real-valued quality score based on some measure of agreement with reference gold output $g$.
-At inference, only input $x$ and generation $y$ are observed; gold output $g$ and metric score $m_{t,k}(y,g)$ are not.
-Instance-level performance prediction. For each task-specific metric $m_{t,k}$, we seek a predictor
+Let $\mathcal{X}$ denote the input space (user queries, upstream agent inputs) and $\mathcal{Y}$ the free-form output space (e.g., text, code, etc.) for possible LLM generations, as well as gold reference outputs, if available.
+We assume each task $t\in\mathcal{T}$ has $K_{t}$ continuous evaluation metrics $\mathcal{M}_{t}=\{m_{t,1},\dots,m_{t,K_{t}}\}$, where each metric $m_{t,k}:\mathcal{Y}\rightarrow\mathbb{R}$ assigns LLM output $y$ a real-valued quality score based on a measure of quality – for example, direct human judgment or agreement with reference gold output $g$.
+At inference, only input $x$ and generation $y$ are observed; metric score $m_{t,k}(y)$ is not.
+Prediction. For each task-specific metric $m_{t,k}$, we seek a predictor defined by:
 $f_{t,k}:(\mathcal{X}\!\times\!\mathcal{Y})\rightarrow\big(\hat{s},[\ell,u]\big)$
-for instance $(x,y)$ that returns: (i) a point estimate $\hat{s}\in\mathbb{R}$ for metric score $m_{t,k}(y,g)$ and (ii) a $(1-\alpha)$ prediction interval $\ell,u]\subseteq\mathbb{R}$ around $\hat{s}$ such that:
-$\Pr\!\big{(}m_{t,k}(y,g)\in[\ell,u]\big{)}\geq 1-\alpha,$
+for instance $(x,y)$ that returns: (i) a point estimate $\hat{s}\in\mathbb{R}$ for metric score $m_{t,k}(y)$ and (ii) a $(1-\alpha)$ prediction interval $\ell,u]\subseteq\mathbb{R}$ around $\hat{s}$ such that:
+$\Pr\!\big(m_{t,k}(y)\in[\ell,u]\big)\geq 1-\alpha,$
 for a specified confidence level (e.g., $\alpha=0.05$).
 Supervision.
-For each task $t$, assume we have a dataset $\mathcal{D}$ containing $N_{t}$ training instances $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N_{t}}$ for each of $t$’s $K_{t}$ metrics, where $s_{i}=m_{t,k}(y_{i},g_{i})$ denotes score $s_{i}$ for metric $m_{t,k}$, as calculated for generation $y_{i}$ with respect to gold reference $g_{i}$.
+For each task $t$, assume we have a dataset $\mathcal{D}$ containing $N_{t}$ training instances $\{(x_{i},y_{i},s_{i})\}_{i=1}^{N_{t}}$ for each of $t$’s $K_{t}$ metrics, where $s_{i}=m_{t,k}(y_{i})$ denotes the gold score for metric $m_{t,k}$, as assessed for generation $y_{i}$.
 This setup is both metric-agnostic (allows any metric) and model-agnostic (makes no use of model parameters or decoding internals).
-Objective: minimize point-estimation error and also produce calibrated intervals (see Section 5.2).
-Figure 1: Illustrating different regression model predictions (solid lines) and confidence intervals (color shading around the lines).
-In this example, a Bayesian Ridge regression model performs best as predictions are neither over-confident (XGBoost) nor under-confident (Gaussian Process).
-4 Baseline Methods
-Having defined the desired performance predictor $f_{t,k}$ (Section 3), we now describe two baseline methods for instantiating such a predictor.
+Objective: minimize point-estimation error and produce calibrated intervals (see Section 5.2).
+4 Prediction Methods
+Having defined the PIE performance predictor $f_{t,k}$ (Section 3), we now describe two such predictors.
 Confidence-based Regression (CE-Reg): Extract unsupervised confidence estimates as features, then learn a regression model to map features to the continuous metric score and a prediction interval.
-Methods are presented in Sections 4.1 and 4.2.
-Reference-Free LLM-as-a-Judge: Prompt an LLM to directly output a point estimate and prediction interval for the target metric, using labeled examples for in-context learning. See Section 4.3.
+Sections 4.1 and 4.2 provide further detail.
+Reference-free LLM-as-a-Judge (RF-LLMaaJ): Prompt an LLM to directly output a point estimate and prediction interval for the target metric, using labeled examples for in-context learning.
+Section 4.3 provides further detail.
 4.1 Confidence Estimation (CE)
-Verbalized confidence (1S and 2S): Prompting an LLM for its confidence that an output is correct (see Section 2.6).
+Verbalized confidence prompts an LLM for its confidence that an output is correct.
 One-step (1S) produces both output and estimated confidence in one invocation, while two-step (2S) first generates the output and then calls the LLM a second time to estimate output confidence.
-While 1S reduces computation, 2S provides cleaner delineation for benchmarking generation vs.
-confidence estimation.
-Note that while 2S permits different LLMs to be used for each step, our experiments use the same LLM for both steps.
+While 1S reduces computation, 2S provides cleaner delineation for benchmarking PIE methods on model fixed outputs.
+See Appendix A.2 for example prompts.
 Consistency-based: We generate $n$ outputs and measure coherence between them via 5 methods: Degree Matrix (Deg; Lin et al.
@@
 4.2 Confidence-Based Regression (CE-Reg)
-Given a dataset $\mathcal{D}=\{(c_{i},s_{i})\}_{i=1}^{N},$ of $N$ training examples, where each $c_{i}$ is the confidence feature (verbalized or consistency-based) for output $y_{i}$ and $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold score, we fit an auxiliary regression model to learn the conditional distribution $s_{i}\sim p(s\mid c_{i},\theta),$ where $s_{i}=m_{t,k}(y_{i},g_{i})$ is the gold metric score.
+Given a dataset $\mathcal{D}=\{(c_{i},s_{i})\}_{i=1}^{N},$ of $N$ training examples, where each $c_{i}$ is the confidence feature (verbalized or consistency-based) for output $y_{i}$ and $s_{i}=m_{t,k}(y_{i})$ is the gold score, we fit an auxiliary regression model to learn the conditional distribution $s_{i}\sim p(s\mid c_{i},\theta),$ for gold metric score $s_{i}=m_{t,k}(y_{i})$.
 From the fitted distribution we extract a point estimate $\hat{s}_{i}$ and a $(1-\alpha)$ prediction interval $\ell_{i},u_{i}\,]$ (using $\alpha=0.05$) via
 $\ell_{i}=F^{-1}\bigl(\tfrac{\alpha}{2}\bigr),\quad u_{i}=F^{-1}\bigl(1-\tfrac{\alpha}{2}\bigr)$
-where $F$ is the model’s predictive cumulative distribution function.
+where $F$ is the model’s predictive cumulative distribution function (CDF).
 Regression Models.
@@
 Gaussian-error models provide mean $\hat{s} &$ variance $\hat{\sigma}^{2}$ directly; ensembles approximate variance via bootstraps/trees.
-CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
-Eccen.
-EigVal.
-LexSim.
-NumSet.
-1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.45 0.43 0.44 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.45 0.45 0.44 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.15 0.14 0.16 LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.25 0.22 0.23 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.24 0.21 0.21 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.08 0.08 0.08 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.31 0.32 0.34 Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.35 0.34 0.35 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.27 0.28 0.27 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.17 0.16 0.17 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.22 0.22 0.20 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.18 0.17 0.17 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.09 0.09 0.10 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.33 0.32 0.31 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.23 0.22 0.22 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.19 0.18 0.18 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.09 0.09 0.09 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.20 0.20 0.20 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.12 0.12 0.12 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.22 0.20 0.19 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.22 0.20 0.19 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.32 0.34 0.36 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.21 0.20 0.20 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.06 0.06 0.06 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.39 0.40 0.38 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.28 0.29 0.27 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.15 0.12 0.12 AVERAGE 0.14 0.15 0.15 0.14 0.16 0.15 0.15 0.23 0.22 0.22
-Table 1: RMSE Benchmark.
-We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings.
-BiGGen-BR = BiGGen-Bench-Results.
-For RMSE, values lower the better.
-Judge(Func.) = CodeJudge Functional Correctness.
-Judge(Incon.) = CodeJudge Inconsistency.
-LLMEval(A.) = LLM Eval Accuracy.
-LLMEval(I.) = LLMEval Informativeness.
-CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
-Eccen.
-EigVal.
-LexSim.
-NumSet.
-1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.36 0.37 0.38 0.37 0.37 0.38 0.38 0.26 0.25 0.26 Judge(Incon.) 0.28 0.31 0.32 0.31 0.32 0.32 0.32 0.33 0.34 0.32 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.08 0.08 0.08 ASQA LLMEval(A.) 0.16 0.16 0.16 0.15 0.17 0.16 0.16 0.14 0.14 0.14 LLMEval(I.) 0.16 0.16 0.16 0.16 0.16 0.15 0.16 0.14 0.13 0.14 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.03 0.03 0.03 0.03 BiGGen Judge 0.24 0.30 0.27 0.26 0.31 0.30 0.30 0.20 0.20 0.22 BigCodeBench Judge(Func.) 0.26 0.26 0.27 0.25 0.28 0.28 0.28 0.23 0.23 0.23 Judge(Incon.) 0.21 0.21 0.22 0.21 0.21 0.21 0.21 0.18 0.19 0.18 BigCodeBench CodeBLEU 0.08 0.09 0.09 0.08 0.10 0.10 0.10 0.10 0.10 0.10 LLMEval(A.) 0.16 0.17 0.16 0.17 0.17 0.17 0.17 0.14 0.14 0.13 LLMEval(I.) 0.13 0.13 0.14 0.14 0.12 0.13 0.12 0.11 0.10 0.10 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.03 Fact-Bio FActScore 0.24 0.33 0.23 0.21 0.33 0.25 0.32 0.23 0.22 0.20 LLMEval(A.) 0.17 0.18 0.17 0.17 0.18 0.18 0.18 0.14 0.14 0.13 LLMEval(I.) 0.15 0.15 0.15 0.15 0.16 0.15 0.15 0.12 0.11 0.11 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.03 Opus-100 GEMBA 0.13 0.14 0.13 0.12 0.13 0.14 0.14 0.11 0.10 0.10 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.07 0.07 0.07 G-Eval(Coh.) 0.13 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 G-Eval(Con.) 0.11 0.11 0.11 0.11 0.11 0.11 0.11 0.13 0.12 0.11 G-Eval(Flu.) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.27 0.29 0.31 G-Eval(Rel.) 0.12 0.13 0.13 0.13 0.13 0.13 0.13 0.14 0.13 0.12 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02 LLMEval(A.) 0.25 0.30 0.26 0.26 0.30 0.30 0.29 0.27 0.28 0.26 LLMEval(I.) 0.22 0.22 0.22 0.22 0.23 0.22 0.23 0.19 0.20 0.18 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.06 0.05 0.05 AVERAGE 0.07 0.11 0.08 0.07 0.12 0.12 0.12 0.14 0.14 0.14
-Table 2: CRPS Benchmark.
-We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings.
-For CRPS, lower is better.
 4.3 Reference-free LLM-as-a-Judge
-Building on prior work (Section 2.6), we define a Reference-free LLM-as-a-Judge (RF-LLMaaJ) baseline for continuous-score performance prediction using in-context learning Vacareanu et al.
+Building on prior work (Section 2.3), we also compare a reference-free LLM-as-a-Judge (RF-LLMaaJ) for continuous-score estimation using in-context learning Vacareanu et al.
 (2024).
-It is prompted with: 1) a brief description of the task and the scoring metric; 2) $N$ demonstration examples in form (input, output, gold metric score) of the desired scoring format; 3) instructions to produce a point score and 95% upper/lower interval bound; and 4) A test pair (input, output) for which it must predict a point estimate and prediction interval for the target metric.
-5 Benchmark
+It is prompted with: 1) a brief description of the task and the scoring metric; 2) $N$ demonstration examples in form (input, output, gold metric score); 3) instructions to produce a point score and 95% upper/lower interval bound; and 4) A test pair (input, output) for which it must predict a point estimate and prediction interval for the target metric.
+Figure 1: Illustrating different regression model predictions (solid lines) and confidence intervals (color shading around the lines).
+In this example, a Bayesian Ridge regression model performs best as predictions are neither over-confident (XGBoost) nor under-confident (Gaussian Process).
+5 Experiments
 5.1 Experimental Setup
@@
 Metrics include: BERTScore (Zhang* et al., 2020), CodeBLEU (Ren et al., 2020), and reference-based (RB) LLM-as-a-judge metrics, such as LLM-Eval (accuracy and informativeness) for QA (Xian et al., 2025; Lin and Chen, 2023) and BiGGen (Kim et al., 2025), GEMBA_ref (with references) (Kocmi and Federmann, 2023) for MT, G-Eval (coherence, consistency, fluency, and relevancy) (Liu et al., 2023) for text summarization, FActScore (Min et al., 2023) for fact-checking, and CodeJudge Tong and Zhang (2024) for code generation and function-calling (functional correctness and inconsistency).
-These LLM judge methods support more robust evaluation by considering diverse aspects of quality (e.g., factual correctness, informativeness, fluency, relevance, etc.).
-We also include a “direct assessment” (DA) task (predicting human-annotated quality ratings instead of automatic metric scores, see Section 2.4), for the BiGGen-Bench-Results dataset (Kim et al., 2025).
-Because this dataset provides human evaluations for its pre-generated LLM outputs, we use both these annotations and outputs for this DA task.
-However, because the dataset only contain one output per rating, this precludes evaluation of consistency-based methods (see Section 4.1).
+These LLM judge methods support more robust evaluation by considering diverse aspects of quality, such as factual correctness, informativeness, fluency, relevance, etc.
+Direct Assessment (DA) Task.
+To show PIE’s generality in supporting prediction of both directly observed quantities as well as derived evaluation metrics, we include a proof-of-concept evaluation of predicting observed human quality ratings for the BiGGen-Bench-Results dataset (Kim et al., 2025) (Section 2.2).
+This dataset provides human ratings for its pre-generated LLM outputs, and we use both its LLM outputs and human ratings.
+However, because the dataset only contains one output per rating, this precludes evaluation of consistency-based methods (see Section 4.1).
 LLMs.
-Llama 3.2 11B (Meta, 2024) is used in both generating outputs and in our baseline predictors (i.e., for consistency and verbalized confidence methods, as well as our RF-LLMaaJ).
+Llama 3.2-11B (Meta, 2024) and Gemini-2.0-Flash-Lite (Team et al., 2023) are used to produce output generations.
+For PIE prediction, whichever LLM was used for generation is then also used for both predictors, CE-Reg and RF-LLMaaJ.
 We set max tokens =769=769, temperature =1=1, and topP =0.9=0.9.
-To avoid risk of self-enhancement bias (Zheng et al., 2023) in evaluation, we use a different LLM, Gemini-2.0-Flash-Lite, for reference-based (RB) LLM-as-a-judge evaluation metrics.
-Our proof-of-concept direct assessment evaluation setting differs by using human quality ratings and existing LLM outputs: Kim et al. (2025)’s GPT-4-0125 (Achiam et al., 2023) generations.
-Baselines.
+To avoid any self-enhancement bias (Zheng et al., 2023) in ensuring high quality evaluation scores to predict, we use Gemini-2.5-Flash (Team et al., 2023) for reference-based (RB) LLMaaJ evaluation metrics.
+The DA task above differs in using existing LLM outputs from prior work: Kim et al. (2025)’s GPT-4-0125 (Achiam et al., 2023) generations.
+PIE Predictors.
 Consistency methods are implemented by LM-Polygraph (Fadeeva et al., 2023) and computed over n=3n\!=\!3 outputs, with all $n$ outputs sharing the same consistency score.
@@
 Beta regression is implemented in-house.
-For regression model selection, we compare RMSE and CRPS results on the dev set.
+For regression model selection, we compare results on the dev set.
 Based on this (Figure 4), we select Random Forest for our test set evaluation.
-Regarding scale of supervision, for RF-LLMaaJ in-context learning, we evaluate varying $N\in\{4,8,16\}$ demonstration examples.
-We only provide up to 16 examples to the RF-LLMaaJ prompt due to high token costs (especially with long-form outputs), and potentially even degrade performance Levy et al.
+For RF-LLMaaJ in-context learning, we vary $N\in\{4,8,16\}$ demonstration examples.
+We cap at 16 demonstration examples due to high token costs (especially with free-form outputs) and potential to degrade performance Levy et al.
 (2024); Modarressi et al.
 (2025).
-In contrast, CE-R is trivial to train at scale.
+In contrast, CE-Reg is far more efficient to train at scale.
 As for relative LLM compute costs of different methods (beyond the output generation call): consistency methods require n1n-1 additional calls (to generate additional outputs for measuring output consistency), verbalized 1S requires no additional calls, verbalized 2S requires one additional call, and RF-LLMaaJ also requires one additional call.
@@
 Appendix A.2 lists prompts and parsing rules.
-5.2 Evaluating Performance Prediction
-$f_{t,k}$ (Sec. 3) must return both a point estimate $\hat{s}$ and a predictive distribution (from which we take a $(1-\alpha)$ interval), and we evaluate these separately.
-Point estimates.
-We report RMSE between prediction $\hat{s}$ and gold score $q$, providing an interpretable measure of average error magnitude in the target’s original units and serving as our primary evaluation metric.
-Pearson Correlation (PC) is shown only as a diagnostic for feature usefulness (Appendix 9); it does not directly reflect absolute error, and correlation may be high while still incurring large systematic errors.
-However, as it is unscaled, it is useful for comparing between tasks and metrics.
-Prediction intervals.
-Expected Calibration Error (ECE) is popular for classification, measuring how well predicted probabilities match observed frequencies via binning (Naeini et al., 2015).
-However, ECE is designed for binary scores, not for continuous ones.
-Prediction intervals for continuous scores can be visualized in Figure 1, showing each instance’s true score on the y-axis as well as the prediction regressor on the x-axis.
-Ideally, predictions are not too narrow (overconfident) or too wide (underconfident), and several metrics exist for assessing such miscalibrations.
-We use the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007) as our primary metric for evaluating prediction intervals.
-It is a proper scoring rule that jointly rewards calibration and sharpness of the full predictive CDF $F$, where lower CRPS is better:
+CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
+Eccen.
+EigVal.
+LexSim.
+NumSet.
+1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.22 0.27 0.28 0.25 0.33 0.36 0.34 0.28 0.27 0.26 Judge(Incon.) 0.05 0.09 0.09 0.08 0.11 0.12 0.11 0.14 0.14 0.13 APIGen CodeBLEU 0.02 0.04 0.04 0.03 0.05 0.05 0.05 0.06 0.06 0.06 ASQA LLMEval(A.) 0.14 0.18 0.13 0.13 0.20 0.18 0.19 0.16 0.16 0.16 LLMEval(I.) 0.09 0.13 0.09 0.09 0.14 0.14 0.14 0.13 0.11 0.13 ASQA BERTScore 0.02 0.02 0.02 0.01 0.02 0.02 0.02 0.03 0.02 0.02 BiGGen Judge 0.17 0.26 0.18 0.17 0.26 0.26 0.24 0.25 0.26 0.25 BigCodeBench Judge(Func.) 0.21 0.24 0.22 0.22 0.29 0.29 0.28 0.35 0.37 0.38 Judge(Incon.) 0.03 0.04 0.04 0.04 0.07 0.07 0.06 0.12 0.10 0.09 BigCodeBench CodeBLEU 0.05 0.06 0.05 0.05 0.07 0.07 0.07 0.09 0.08 0.08 LLMEval(A.) 0.15 0.25 0.15 0.15 0.24 0.25 0.23 0.23 0.24 0.23 LLMEval(I.) 0.10 0.14 0.10 0.10 0.13 0.14 0.13 0.14 0.13 0.13 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.02 Fact-Bio FActScore 0.10 0.26 0.10 0.11 0.26 0.19 0.26 0.23 0.21 0.22 LLMEval(A.) 0.16 0.25 0.16 0.16 0.25 0.25 0.25 0.23 0.23 0.22 LLMEval(I.) 0.11 0.13 0.11 0.11 0.14 0.14 0.13 0.14 0.13 0.13 MED LF QA BERTScore 0.02 0.02 0.02 0.01 0.02 0.02 0.02 0.03 0.03 0.03 Opus-100 GEMBA 0.10 0.14 0.11 0.11 0.17 0.16 0.16 0.16 0.14 0.14 Opus-100 BERTScore 0.03 0.05 0.04 0.03 0.06 0.06 0.06 0.07 0.07 0.07 G-Eval(Coh.) 0.04 0.05 0.04 0.04 0.05 0.05 0.05 0.08 0.07 0.08 G-Eval(Con.) 0.06 0.08 0.05 0.06 0.08 0.08 0.07 0.10 0.09 0.10 G-Eval(Flu.) 0.01 0.02 0.01 0.01 0.02 0.02 0.02 0.03 0.03 0.03 G-Eval(Rel.) 0.11 0.14 0.11 0.11 0.14 0.14 0.14 0.14 0.14 0.14 Summeval BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.02 0.02 0.02 LLMEval(A.) 0.16 0.29 0.16 0.16 0.31 0.29 0.28 0.30 0.30 0.30 LLMEval(I.) 0.16 0.23 0.17 0.16 0.25 0.24 0.23 0.25 0.25 0.24 TruthfulQA BERTScore 0.02 0.03 0.02 0.02 0.03 0.03 0.03 0.06 0.05 0.05 AVERAGE 0.09 0.13 0.09 0.09 0.14 0.13 0.13 0.14 0.14 0.14
+Table 3: CRPS Benchmark.
+We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings.
+CE-Reg generally scores better (lower) CRPS than RF-LLMaaJ.
+5.2 PIE Evaluation
+Predictor $f_{t,k}$ (Section 3) must return both a point estimate $\hat{s}$ and a predictive distribution (with a $(1-\alpha)$ interval).
+We evaluate these with 4 PIE metrics (not to be confused with task-specific metrics).
+Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery, 2007) is PIE’s primary metric for evaluating both point estimates and prediction intervals.
+It is a proper scoring rule that jointly rewards calibration and sharpness of the full predictive CDF $F$; lower CRPS is better:
 $\mathrm{CRPS}(F,s)=\int_{-\infty}^{\infty}\big(F(x)-\mathbf{1}\{x\geq q\}\big)^{2}\,dx$
-For interpretability, we also measure Average Coverage Error (ACE) (Shrivastava et al., 2016; Xie et al., 2023), which compares empirical coverage of the nominal $(1-\alpha)$ interval to its target as $\mathrm{ACE}=\bigl{|}\mathrm{PICP}(p)-p\bigr{|}$ where
+Prediction intervals for continuous scores can be visualized in Figure 1, showing each instance’s true score on the y-axis and the prediction regressor on the x-axis.
+Ideally, predictions are neither too narrow (overconfident) nor too wide (underconfident).
+While Expected Calibration Error (ECE) is widely used for classification to compare predicted probabilities with observed frequencies via binning (Naeini et al., 2015), it is traditionally defined over binary scores, not continuous ones.
+Kuleshov et al.
+(2018) propose an ECE variation for continuous scores based on quantile coverage, but unlike CRPS, such diagnostics do not reward distribution sharpness and can therefore favor overly conservative (wide) intervals (Gneiting et al., 2007).
+Other complementary metrics We also report RMSE (Appendix 8), providing an interpretable measure of average error magnitude in the target’s original units.
+Pearson Correlation (PC) is also used as a diagnostic for feature usefulness (Appendix 9); it does not directly reflect absolute error, and correlation may be high while still incurring large systematic errors.
+However, as it is unscaled, it is useful for comparing between tasks and metrics and for reflecting instance-level discriminative predictability.
+In addition, we also measure Average Coverage Error (ACE) (Shrivastava et al., 2016; Xie et al., 2023) (Appendix 10), which compares empirical coverage of the nominal $(1-\alpha)$ interval to its target as $\mathrm{ACE}=\bigl|\mathrm{PICP}(p)-p\bigr|$ where
 $\mathrm{PICP}(p)$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\,\ell_{i}(p)\leq s_{i}\leq u_{i}(p)\,\}$
 and $p=1-\alpha$ (we use $p=0.95$).
-Lower ACE indicates better calibration, $p>0$ indicates the prediction is under-confident, and $p<0$ indicates it is over-confident.
+Lower ACE indicates better calibration, with positive $p$ reflecting under-confidence and negative $p$ reflecting over-confidence.
 Note that one disadvantage of ACE is that it does not penalize excessively wide intervals (e.g., under-confident intervals that always cover $s_{i}$ incur a maximum $0.05$ penalty when $p=0.95$) as much as narrow ones.
-Hence, we treat ACE as a secondary, human-readable diagnostic and rely on CRPS for model selection.
-6 Results
-In this section, we address three central questions: 1) Is instance-level performance prediction of continuous, reference-based metrics for long-form LLM outputs feasible?
-2) How much supervision is necessary to achieve reasonable performance?
-and 3) How do confidence-based regression and reference-free LLM-as-a-Judge (Section 4) compare?
-We evaluate all baselines across 11 datasets and multiple metrics (Section 5.1) using RMSE to judge point estimate accuracy and CRPS to judge calibration of predictive distributions, with ACE and PC used for additional diagnostics (Section 5.2).
-Our key findings are as follows.
-Performance prediction is indeed feasible: regression on off-the-shelf confidence estimates achieves an average RMSE of 0.14 and CRPS of 0.08, with intervals calibrated to within 16% of nominal coverage and requiring as few as 16 training examples to converge.
-Tables 1 and 2 show that regression on the consistency-based “DegMat” confidence estimate performs best in terms of both RMSE and CRPS.
-Only 16 training instances are required for the CE-Reg’s test loss to stabilize.
-As long-form gold references can be expensive to collect, it is important to understand how sample-efficient these performance prediction methods would be.
+Hence, we treat ACE as a secondary, human-readable diagnostic and rely primarily on CRPS for PIE evaluation.
+Figure 2: CE-Reg outperforms RF-LLMaaJ on RMSE, CRPS, Correlation, and ACE (95%) on Llama 3.2 11B (left) and Gemini 2.0 flash-lite (right), averaged across all datasets and task metrics.
+Lower is better for all except correlation.
+This result shows minimal difference between LLMs.
+6 Findings
+We now summarize 6 key empirical findings from evaluating PIE over 11 free-form generation tasks, 15 metrics, and 2 backbone LLMs.
+Rather than focus solely on point estimates or preference judgments, our experiments jointly assess the prediction of continuous evaluation scores and their associated uncertainty.
+We organize the results to study key properties of PIE, including feasibility, sample efficiency, metric-dependent behavior, and robustness across estimation techniques and models.
+PIE is feasible Across datasets and target metrics, CE-based regression achieves an average CRPS of 0.10, RMSE of 0.16, and PC of 0.34 (Tables 3, 8, and 9), indicating that simple confidence features support useful prediction of continuous quality scores and calibrated uncertainty at inference time.
+This holds both for derived evaluation metrics and for human ratings (Table 4).
+Performance also stabilizes with as few as 16 labeled examples, suggesting PIE can be practical when annotation budgets are small.
+PIE supports sample-efficient predictors As free-form gold references can be expensive to collect, it is important to understand how sample-efficient these methods would be.
 To answer that, we vary the size of the training set from as few as 4 instances and doubling up to the full training data size for CE-Reg.
-For RF-LLMaaJ we limit up to 16 instances.
-CE-Reg achieves close to our best observed accuracy with remarkably few labeled examples: Figure 2 shows that RMSE converges after just 16 training instances across all datasets.
-Even when matched to RF-LLMaaJ’s maximum of 16 in-context examples (Table 4), CE-Reg delivers substantially lower error.
-Although performance varies slightly across datasets, the convergence point is consistent, demonstrating that this regression approach to performance prediction can be effectively trained with a modest sample size, making it easy to collect training data even for out-of-domain tasks and mitigating potential domain differences.
-CE-Reg outperforms RF-LLMaaJ.
-In Tables 1 and 2, CE-Reg methods achieve RMSEs averaging in the 0.14–0.15 range and reaching 0.35 at worst, depending on the task metric.
-Meanwhile, RF-LLMaaJ incurs much larger errors (averaging around 0.29, 0.85 at worst).
-Similarly, CE-Reg attains CRPS values between 0.07–0.12 averaged over task metrics, compared to about 0.21 for RF-LLMaaJ.
-Interestingly, most methods exhibit an over-confidence (intervals are too narrow) in ACE (95%) (Appendix 8), with verbalized CE-Reg producing ACEs of –0.89 and –0.79 and RF-LLMaaJ yielding ACEs of –0.51 and –0.52, while only some of the consistency-based features come close to nominal coverage (e.g.
-DegMat ACE: –0.16).
-Notably, this pattern largely reflects our use of RandomForest regression: substituting a XGBoost model markedly improves ACE, albeit at the expense of other performance metrics (see Figure 3 in Appendix).
-In addition, Judge scores and human ratings in the BiGGen-Bench-Results dataset yield similar performance, indicating our baselines generalize beyond evaluation metrics.
-Prediction error is primarily metric-dependent.
-Across most tasks and methods, CodeBLEU and BERTScore are the easiest metrics to predict: CE-Reg consistently achieves low RMSE and CRPS and high PC on these metrics, and RF-LLMaaJ exhibits a similar trend on most datasets.
-The G-Eval in text summarization tasks are consistently low in RMSE and CRPS but also low in PC – this can be explained by the fact that its target score distribution is tightly concentrated.
-CodeJudge and LLMEvals tend to be the most difficult with high RMSE and CRPS and relatively low PC.
-Variation in RMSE and CRPS across evaluation metrics shows that prediction difficulty is largely metric dependent, with low metric-to-metric variability within each task.
-In our findings, we do not observe a similar trend for task-to-task variability.
-Therefore, we do not conclude anything about relative inherent difficulty between tasks.
-Graph-based consistency features consistently outperform all other CE-Reg methods across every dataset and metric Among all CE-Reg, the four graph-based consistency measures (DegMat, Eccentricity, EigVal, LexSim) achieve the lowest CRPS, indicating that capturing variability across multiple generations provides a stronger predictive signal than one- or two-step verbalized confidence prompts, on all types of dataset or metrics.
+For RF-LLMaaJ, we limit up to 16 instances.
+CE-Reg achieves close to our best observed accuracy using few labeled examples: Figure 3 shows that CRPS converges after just 16 training instances across all datasets.
+Even when compared to RF-LLMaaJ’s maximum of 16 in-context examples (Table 5), CE-Reg delivers substantially lower error.
+Although performance varies slightly across datasets, the convergence point is consistent.
+This shows that the regression approach can be effectively trained with a small sample size, making it easy to collect training data even for out-of-domain tasks to mitigate potential domain differences.
+PIE is more accurately addressed by CE-Reg than by RF-LLMaaJ.
+Because PIE requires predicting continuous evaluation metrics together with uncertainty at the instance level, not all evaluation paradigms are equally well matched to this objective.
+We therefore compare CE-Reg and RF-LLMaaJ as two representative paradigms for performance prediction under PIE.
+Across datasets and evaluation metrics, CE-Reg consistently achieves substantially lower error under the primary CRPS objective than RF-LLMaaJ (Table 3).
+Averaged over task metrics, CE-Reg attains CRPS values averaging 0.11, compared to 0.14 for RF-LLMaaJ, indicating more informative predictive distributions.
+This advantage is also reflected in point accuracy: CE-Reg achieves RMSEs averaging 0.16, whereas RF-LLMaaJ incurs much larger errors, averaging around 0.22 (Table 8).
+In addition, CE-Reg preserves stronger instance-level ordering, achieving consistently higher correlation of 0.28 across metrics than RF-LLMaaJ’s 0.09 (Table 9).
+Notably, for many task–metrics, RF-LLMaaJ exhibits near-zero correlation, indicating limited ability to recover instance-level signal.
+Increasing in-context examples from 4-16 yields negligible improvement (Table 5), suggesting that additional prompting alone does not reliably address this.
+One plausible explanation is RF-LLMaaJ’s sensitivity to prompt formulation and target specification.
+Although both RF-LLMaaJ and verbalized confidence rely on in-context LLM predictions, verbalized confidence appears more stable across tasks, whereas RF-LLMaaJ must predict diverse task-specific metrics, which few-shot prompting may not reliably support.
+Taken together, these findings indicate that confidence estimates are informative signals for PIE but require supervision for effective calibration, while RF-LLMaaJ approaches remain less reliable across metrics and evaluation sources.
+PIE benefits most from graph-based consistency features.
+Among all CE-Reg, the four graph-based consistency measures (DegMat, Eccentricity, EigVal, LexSim) achieve the lowest CRPS, indicating that capturing variability across multiple generations provides a stronger predictive signal than one- or two-step verbalized confidence prompts, on all types of datasets and metrics.
 By contrast, the NumSemSets consistency metric performs on par with the verbalized methods, trailing the graph-based techniques.
 Better results from graph-based consistency over NumSemSets and verbalized confidence are more obviously reflected in the ACE and PC measures (Tables 10 and 9).
-RF-LLMaaJ is a poor performance predictor.
-Both point prediction and interval estimation are ill-calibrated compared to CE-Reg methods.
-In fact, only on a few tasks does RF-LLMaaJ seem to have any ability to predict evaluation metrics at all – Appendix Table 9 shows near-zero correlation between predicted and actual scores for the majority of task metrics.
-Furthermore, increasing in-context examples from 4 to 16 has negligible benefit.
-There are some task metrics where RF-LLMaaJ significantly better than random chance, but the variability across tasks and metrics highlights how sensitive and unreliable this approach can be, relative to the consistent performance seen in the consistency-based confidence regression approach.
-One reason that RF-LLMaaJ struggles to predict evaluation metrics could be its sensitivity to the prompt.
-A key insight is that, despite both using in-prompt LLM predictions, verbalized confidence is overall more reliable than RF-LLMaaJ, which is prompted to predict arbitrary task metrics.
-This may reflect the LLM’s pre-training data, which likely contains the general notion of confidence much more so than task-specific metrics like CodeBLEU or GEMBA; and the few-shot examples are not enough to overcome this deficiency.
+PIE yields consistent results across LLMs.
+Figure 2 evaluates the generality of CE-Reg and RF-LLMaaJ by comparing their performance across two backbone models.
+In addition to Llama 3.2 11B, we re-ran the same experimental setup on Gemini 2.0 Flash-Lite.
+Across both models and all evaluation criteria, we observe that CE-Reg outperforms RF-LLMaaJ, with absolute performance levels also comparable across models.
+This consistency suggests that the advantages of CE-Reg are not model-specific and that PIE generalizes across LLM families.
+PIE performance varies across task metrics.
+Across datasets, BERTScore and CodeBLEU achieve low CRPS (Table 3) and high correlation (Table 9), indicating that their per-instance predictive distributions are accurate in absolute magnitude and that predicted scores are well-aligned in relative terms across instances.
+G-Eval metrics likewise attain low CRPS, but their correlation is consistently low, suggesting a different regime: predictions match typical score levels (often because true scores are tightly concentrated), yet are less informative for distinguishing instances.
+In contrast, LLMEval, BiGGen Judge, GEMBA, and FactScore produce broader target distributions, leading to slightly higher CRPS while still maintaining relatively high correlation.
+This pattern suggests that these metrics contain more instance-to-instance variation in quality, enabling meaningful relative comparisons even when uncertainty-aware prediction is harder.
+Overall, different tasks and metrics exhibit different mixtures of (i) distributional predictability of absolute scores and (ii) relative, instance-discriminative predictability.
 CE-Reg: Verbalized RF-LLMaaJ
 2S 4-shot 8-shot 16-shot
-RMSE 0.26 0.32 0.32 0.34
 CRPS 0.21 0.22 0.22 0.24
 Table 4: Direct Assessment: For BiGGen we evaluate two baselines using pre-generated outputs—Verbalized 2S and RF-LLMaaJ—for directly predicting human ratings.
-Consistency-based (requiring additional outputs) and the Verbalized 1S (requiring prompt re-design) are omitted.
+Methods requiring additional outputs are omitted.
+This result shows CE-Reg calibrating better than LLMaaJ on predicting human judgments.
 CE-Reg: DegMat RF-LLMaaJ
 4 inst. 8 inst. 16 inst. 4-shot 8-shot 16-shot
-RMSE 0.17 0.18 0.17 0.23 0.22 0.22
-CRPS 0.10 0.10 0.10 0.14 0.14 0.14
-Table 4: Comparison of RMSE and CRPS for CE-Reg using DegMat feature versus reference-free LLM-as-a-Judge (RF-LLMaaJ) using different number of training examples.
-(a) RMSE vs. Training Data Instances
-(b) CRPS vs. Training Data Instances
-Figure 2: Impact of Training Data Size on CE-Reg Performance and Test Set RMSE and CRPS Across Datasets, using the selected task-specific metric (see Table 6).
-The plot shows RMSE and CRPS values for training data sizes capped at 64 instances.
-Notably, both stabilize after 16 instances, suggesting a performance plateau for certain datasets
+CRPS 0.09 0.09 0.09 0.14 0.14 0.14
+Table 5: Comparing CE-Reg using DegMat feature vs. RF-LLMaaJ, varying training example count, averaged across all datasets, showing undetectable benefit from additional training examples on CRPS.
+Figure 3: Impact of training data size on CE-Reg performance, measured by test-set CRPS across datasets using the selected task-specific metric (see Table 7).
+Training data sizes are capped at 64 instances.
+CRPS stabilizes after approximately 16 instances, suggesting a plateau for certain datasets.
 7 Conclusion
-This work introduces the task of instance-level performance prediction for long-form LLM generation, in which a black-box model’s input–output pair is used to predict arbitrary continuous, reference-based metric scores with associated prediction intervals.
-To support progress, we release a comprehensive benchmark spanning 11 datasets across summarization, machine translation, code generation, and question-answering, with multiple task evaluation metrics to predict.
-Our experiments compare two families of performance prediction approaches as baselines.
-The first–confidence-based regression–extracts off-the-shelf verbalized and consistency-based confidence signals, then trains classic regression models to output both point estimates and prediction intervals.
-The second–a reference-free LLM-as-a-Judge–asks an LLM, via in-context prompting, to yield metric estimates and 95% intervals for each input-output-instance.
-We find that confidence-based regression, particularly when using consistency features, yields substantially better RMSE and CRPS than our implemented LLM-based judging.
-This demonstrates that lightweight supervised regression of simple confidence estimates can outperform few-shot LLM judging for both accuracy and uncertainty quantification.
-Our work lays the groundwork for research in uncertainty-aware, per-instance continuous-scored performance prediction.
-We hope our benchmark, code, and baselines will encourage community efforts toward robust, uncertainty-aware evaluation and deployment of long-form LLMs.
+In this work, we propose the task of Performance Interval Estimation (PIE) for predicting quality of free-form LLM generations.
+While confidence estimation is well-known for predicting the probability for whether model outputs are correct or not, predicting binary correctness in this way is most sensible when generation tasks have exact answers.
+In contrast, free-form generations are often more nuanced, with output quality being both fine-grained and multi-faceted.
+We therefore motivate prediction of arbitrary continuous, reference-based metric scores (point estimates) with calibrated prediction intervals.
+Framing this prediction task as regression allows us to be agnostic as to both: 1) whether the prediction target is ordinal or continuous; and 2) whether we are predicting an observed value (e.g., human rating or program execution time) or a derived evaluation metric.
+We emphasize the importance of prediction intervals for risk-informed decision-making; a confidence interval lower-bounds expected quality to better assess the level of risk.
+To promote this prediction task and foster community progress, we release our dataset, PIE-data, spanning 11 datasets across summarization, machine translation, code generation, function-calling, and question-answering, with multiple task evaluation metrics to predict.
+We also share our implemented methods and their outputs.
+Our experiments compare two alternative prediction methods: confidence-based regression (CE-Reg), which extracts off-the-shelf verbalized and consistency-based confidence signals and trains classic regression models to output both point estimates and prediction intervals; and reference-free LLM-as-a-judge (RF-LLMaaJ), which prompts an LLM with in-context learning to yield metric estimates and 95% intervals for each input-output-instance.
+CE-Reg is substantially more accurate and better calibrated than RF-LLMaaJ, especially with consistency features, showing that lightweight supervised regression over simple confidence signals can outperform few-shot LLM judging for both prediction and uncertainty quantification.
 Acknowledgments
-This research was supported in part by Cisco, Amazon, and by Good Systems [^1] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies).
+This research was supported in part by Cisco, Amazon, and by Good Systems [^3] https://goodsystems.utexas.edu/ (a UT Austin Grand Challenge dedicated to developing responsible AI technologies), as well as CosmicAI [^4] The NSF-Simons AI Institute for Cosmic Origins (CosmicAI): https://www.cosmicai.org/ support from the NSF (Cooperative Agreement 2421782) and the Simons Foundation (grant MPS-AI-00010515).
 We thank the Texas Advanced Computing Center (TACC) for use of its superb infrastructure.
@@
 Read the news article carefully and identify the main facts and key points.
-2.
+{internallinenumbers*} 2.
 Read the summary and compare it to the article.
@@
 Assign a score for {metric} based on the Evaluation Criteria.
-Source Text:        {source}        Summary:        {prediction}            Evaluation Form (scores ONLY), provide your response in the following format:    - {metric}:        evaluation_criteria = {        "Coherence": """Coherence (1-5) - the collective quality of all sentences.
-A well-structured summary should build from sentence to a coherent body of information about a topic.""",        "Consistency": """Consistency (1-5) - the factual alignment between the summary and the        source document.
-A factually consistent summary contains only statements entailed by the source document and avoids hallucinated facts.""",        "Fluency": """Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.
+Source Text:      {source}      Summary:      {prediction}        Evaluation Form (scores ONLY), provide your response in the following format:    - {metric}:      evaluation_criteria = {    {internallinenumbers*} "Coherence": """Coherence (1-5) - the collective quality of all sentences.
+A well-structured summary should build from sentence to a coherent body of information about a topic.""",    "Consistency": """Consistency (1-5) - the factual alignment between the summary and the    {internallinenumbers*} source document.
+A factually consistent summary contains only statements entailed by the source document and avoids hallucinated facts.""",    {internallinenumbers*} "Fluency": """Fluency (1-3): the quality of the summary in terms of grammar, spelling, punctuation, word choice, and sentence structure.
 - 1: Poor.
 The summary has many errors that make it hard to understand or sound unnatural.
-- 2: Fair.
+{internallinenumbers*} - 2: Fair.
 The summary has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible.
 - 3: Good.
-The summary has few or no errors and is easy to read and follow.""",        "Relevance": """Relevance (1-5) - selection of important content from the source.
+The summary has few or no errors and is easy to read and follow.""",    {internallinenumbers*} "Relevance": """Relevance (1-5) - selection of important content from the source.
 The summary should include only important information from the source document.
@@
 Prompts for LLM-based Evaluation Metrics (LLM-Eval for QA) (Xian et al., 2025; Lin and Chen, 2023)
-      Task: Score the following LLM output of a question-answering task with respect to the following aspects using a 1 to 5 star rating system.
-Dataset: The dataset is a Question-Answering dataset, specifically designed for evaluating factual precision and detailed comparative reasoning in AI-generated answers.
-Output: Begin your evaluation by providing a short explanation.
+      {internallinenumbers*} Task: Score the following LLM output of a question-answering task with respect to the following aspects using a 1 to 5 star rating system.
+{internallinenumbers*} Dataset: The dataset is a Question-Answering dataset, specifically designed for evaluating factual precision and detailed comparative reasoning in AI-generated answers.
+{internallinenumbers*} Output: Begin your evaluation by providing a short explanation.
 Be as objective as possible.
@@
 System Prompt:
-      You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
+      {internallinenumbers*} You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
 User Prompt:
-      ###Task Description:    An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing an evaluation criteria are given.
-1.
+      ###Task Description:    {internallinenumbers*} An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing an evaluation criteria are given.
+{internallinenumbers*} 1.
 Write a detailed feedback that assesses the quality of the response strictly based on the given score rubric, not evaluating in general.
-2.
+{internallinenumbers*} 2.
 After writing a feedback, write a score that is an integer between 1 and 5.
 You should refer to the score rubric.
-3.
+{internallinenumbers*} 3.
 The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between 1 and 5)"    4.
@@
 Inconsistency Prompt:
-      You will be provided with a problem statement, a code snippet that supposedly addresses the problem,    and a catalog of code inconsistencies.
+      {internallinenumbers*} You will be provided with a problem statement, a code snippet that supposedly addresses the problem,    and a catalog of code inconsistencies.
 Evaluation Steps:    1.
 Read the problem statement carefully to identify the functionalities required for the    implementation.
-2.
+{internallinenumbers*} 2.
 Read the code snippet and compare it to the problem statement.
@@
 Output your answer in a JSON format list.
-a) If the code snippet is correct, output: ["inconsistency": "None", "severity": "Negligible"].
-b) If the code snippet is incorrect, output the identified inconsistencies and their severity    according to the catalog of code inconsistencies.
-For example: ["inconsistency": "<inconsistency1>",    "severity": "<severity1>", "inconsistency": "<inconsistency2>", "severity": "<severity2>", ...]    Problem: {PROBLEM}    Code Snippet: {CODE}    Taxonomy of Common Inconsistencies:    1.
+{internallinenumbers*} a) If the code snippet is correct, output: ["inconsistency": "None", "severity": "Negligible"].
+b) If the code snippet is incorrect, output the identified inconsistencies and their severity    {internallinenumbers*} according to the catalog of code inconsistencies.
+For example: ["inconsistency": "<inconsistency1>",    {internallinenumbers*} "severity": "<severity1>", "inconsistency": "<inconsistency2>", "severity": "<severity2>", ...]    Problem: {PROBLEM}    Code Snippet: {CODE}    Taxonomy of Common Inconsistencies:    1.
 Missing dependency declarations: Negligible    2.
@@
 Dataset Type Data Split Size Task-Specific Metrics
-APIGen (Liu et al., 2025) Function Calling 8955 / 2985 / 2985 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
-ASQA (Stelmakh et al., 2022) Question Answering 1689 / 563 / 564 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite], BERTScore
-BigCodeBench (Zhuo et al., 2025) Code Generation 1994 / 665 / 665 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.0-flash-lite]
-BiGGen (Kim et al., 2025) Open-Ended Generation 847 / 283 / 283 LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
+APIGen (Liu et al., 2025) Function Calling 8955 / 2985 / 2985 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.5-flash]
+ASQA (Stelmakh et al., 2022) Question Answering 1689 / 563 / 564 LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash], BERTScore
+BigCodeBench (Zhuo et al., 2025) Code Generation 1994 / 665 / 665 CodeBLEU, CodeJudge (functional correctness, inconsistency)[gemini-2.5-flash]
+BiGGen (Kim et al., 2025) Open-Ended Generation 847 / 283 / 283 BiGGen-Judge [gemini-2.5-flash]
 BiGGen-Bench-Results (Kim et al., 2025) Open-Ended Generation 1662 / 555 / 555 Human Rating
-Eli5 (Fan et al., 2019) Question Answering 2657 / 886 / 886 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
-FActScore-Bio (Min et al., 2023) Fact-Checking 1216 / 406 / 406 FActScore [gpt-3.5-turbo]
-MedLFQA (Hosseini et al., ) Medical QA 8855 / 2952 / 2952 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
-Opus-100 (EN→ES) (Zhang et al., 2020; Tiedemann, 2012) Machine Translation 3579 / 1194 / 1194 BERTScore, GEMBA_SQM_ref [gemini-2.0-flash-lite]
-Summeval (Fabbri et al., 2021) Text Summarization 2809 / 937 / 937 BERTScore, G-Eval (Coherence, Consistency, Fluency, Relevance) [gemini-2.0-flash-lite]
-TruthfulQA (Lin et al., 2022) Question Answering 1455 / 486 / 486 BERTScore, LLM-EVAL (accuracy, informativeness) [gemini-2.0-flash-lite]
+Eli5 (Fan et al., 2019) Question Answering 2657 / 886 / 886 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
+Fact-Bio (Min et al., 2023) Fact-Checking 1216 / 406 / 406 FActScore [gpt-3.5-turbo]
+Med LF QA (Hosseini et al., ) Medical QA 8855 / 2952 / 2952 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
+Opus-100 (EN→ES) (Zhang et al., 2020; Tiedemann, 2012) Machine Translation 3579 / 1194 / 1194 BERTScore, GEMBA_SQM_ref [gemini-2.5-flash]
+Summeval (Fabbri et al., 2021) Text Summarization 2809 / 937 / 937 BERTScore, G-Eval (Coherence, Consistency, Fluency, Relevance) [gemini-2.5-flash]
+TruthfulQA (Lin et al., 2022) Question Answering 1455 / 486 / 486 BERTScore, LLM‑EVAL (accuracy, informativeness) [gemini-2.5-flash]
 Table 6: Our study and benchmark cover twelve datasets spanning a range of diverse tasks: code generation, fact-checking, function-calling, machine translation (MT), question-answering (QA), summarization, and others.
@@
 For each dataset, we also specify one or more task-specific, multi-dimensional and partial-credit evaluation metric used to evaluate generation performance.
+BiGGen-Bench-Results is the direct assessment evaluation setting.
 Dataset Selected Metrics
-APIGEN (Liu et al., 2025) CodeBLEU
-ASQA (Stelmakh et al., 2022) BERTScore
-BigCodeBench (Zhuo et al., 2025) CodeBLEU
+APIGen (Liu et al., 2025) Judge (Funct.)
+ASQA (Stelmakh et al., 2022) LLMEval(A.)
+BigCodeBench (Zhuo et al., 2025) Judge (Funct.)
 BiGGen (Kim et al., 2025) Judge
-Eli5 (Fan et al., 2019) BERTScore
-FActscore-Bio (Min et al., 2023) FActScore
-MED LF QA (Hosseini et al., ) BERTScore
-Opus-100 (EN to ES) (Zhang et al., 2020; Tiedemann, 2012) BERTScore
-Summeval (Fabbri et al., 2021) BERTScore
-TruthfulQA (Lin et al., 2022) BERTScore
+Eli5 (Fan et al., 2019) LLMEval(A.)
+Fact-Bio (Min et al., 2023) FActScore
+MED LF QA (Hosseini et al., ) LLMEval(A.)
+Opus-100 (Zhang et al., 2020; Tiedemann, 2012) GEMBA
+Summeval (Fabbri et al., 2021) G-Eval(Rel.)
+TruthfulQA (Lin et al., 2022) LLMEval(A.)
 Table 7: While Table 6 includes our full set of task-specific evaluation metrics, for brevity we sometimes select only a single metric per dataset to report, as listed in this table here.
@@
 A.4 Complementary Tables and Figures
-Dataset Task-Specific Metric Score
-APIGen CodeBLEU 0.32
-Judge(Func.) 0.77
-Judge(Incon.) 0.56
-ASQA BERTScore 0.89
-LLMEval(A.) 0.87
-LLMEval(I.) 0.84
-BiGGen Judge 0.67
-BigCodeBench CodeBLEU 0.11
-Judge(Func.) 0.77
-Judge(Incon.) 0.63
-Summeval G-Eval(Coh.) 0.88
-G-Eval(Con.) 0.95
-G-Eval(Flu.) 0.59
-G-Eval(Rel.) 0.79
-BERTScore 0.90
-Eli5 BERTScore 0.85
-LLMEval(A.) 0.72
-LLMEval(I.) 0.63
-FActScore-Bio FActScore 0.43
-Med-LFQA BERTScore 0.84
-LLMEval(A.) 0.76
-LLMEval(I.) 0.59
-Opus-100 GEMBA-SQM_ref 0.87
-BERTScore 0.85
-TruthfulQA BERTScore 0.87
-LLMEval(A.) 0.60
-LLMEval(I.) 0.58
-Table 7: Llama 3.2 11B performance on 10 long-form datasets with various task-specific evaluation metrics.
-All scores are normalized by their respective scale to [0,1].
-Performance prediction then seeks to predict these gold-standard evaluation scores.
 (a) RMSE for regression models
@@
 Figure 4: Comparison across regression models on development sets. For CRPS, RMSE, ACE(95%) lower is better; for Pearson correlation higher is better.
+CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
+Eccen.
+EigVal.
+LexSim.
+NumSet.
+1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.39 0.38 0.40 0.39 0.37 0.39 0.39 0.43 0.43 0.43 Judge(Incon.) 0.11 0.12 0.12 0.12 0.12 0.12 0.12 0.23 0.21 0.20 APIGen CodeBLEU 0.05 0.06 0.06 0.06 0.06 0.06 0.06 0.12 0.11 0.10 ASQA LLMEval(A.) 0.25 0.24 0.25 0.25 0.23 0.23 0.24 0.26 0.26 0.25 LLMEval(I.) 0.17 0.17 0.17 0.17 0.17 0.17 0.18 0.20 0.18 0.20 ASQA BERTScore 0.02 0.02 0.02 0.02 0.03 0.02 0.02 0.08 0.05 0.05 BiGGen Judge 0.30 0.31 0.29 0.28 0.30 0.30 0.31 0.35 0.36 0.35 BigCodeBench Judge(Func.) 0.36 0.34 0.36 0.37 0.35 0.35 0.35 0.48 0.49 0.50 Judge(Incon.) 0.09 0.09 0.09 0.09 0.09 0.09 0.09 0.22 0.19 0.17 BigCodeBench CodeBLEU 0.08 0.09 0.08 0.08 0.10 0.10 0.10 0.14 0.13 0.13 LLMEval(A.) 0.26 0.28 0.26 0.26 0.27 0.28 0.28 0.33 0.33 0.32 LLMEval(I.) 0.17 0.18 0.16 0.17 0.18 0.18 0.18 0.21 0.21 0.21 Eli5 BERTScore 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.06 0.07 0.05 Fact-Bio FActScore 0.20 0.31 0.20 0.22 0.31 0.25 0.32 0.33 0.31 0.34 LLMEval(A.) 0.28 0.29 0.28 0.28 0.28 0.28 0.29 0.34 0.34 0.32 LLMEval(I.) 0.18 0.19 0.18 0.18 0.19 0.18 0.18 0.21 0.21 0.21 MED LF QA BERTScore 0.03 0.03 0.03 0.03 0.03 0.03 0.03 0.09 0.09 0.09 Opus-100 GEMBA 0.21 0.22 0.22 0.21 0.22 0.22 0.23 0.26 0.24 0.24 Opus-100 BERTScore 0.06 0.07 0.07 0.06 0.08 0.08 0.08 0.12 0.12 0.12 G-Eval(Coh.) 0.14 0.12 0.13 0.14 0.12 0.12 0.12 0.16 0.16 0.16 G-Eval(Con.) 0.18 0.15 0.17 0.17 0.14 0.15 0.15 0.19 0.18 0.20 G-Eval(Flu.) 0.05 0.05 0.05 0.06 0.05 0.05 0.05 0.07 0.07 0.07 G-Eval(Rel.) 0.20 0.18 0.20 0.21 0.18 0.18 0.18 0.21 0.22 0.22 Summeval BERTScore 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.06 0.06 0.06 LLMEval(A.) 0.31 0.34 0.30 0.30 0.35 0.34 0.34 0.44 0.43 0.43 LLMEval(I.) 0.27 0.28 0.27 0.26 0.29 0.29 0.29 0.35 0.36 0.34 TruthfulQA BERTScore 0.04 0.04 0.04 0.04 0.04 0.04 0.04 0.15 0.12 0.12 AVERAGE 0.16 0.17 0.16 0.16 0.17 0.17 0.17 0.23 0.22 0.22
+Table 8: RMSE Benchmark.
+We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings.
+For RMSE, lower is better.
+Judge(Func.) = CodeJudge Functional Correctness.
+Judge(Incon.) = CodeJudge Inconsistency.
+LLMEval(A.) = LLM Eval Accuracy.
+LLMEval(I.) = LLMEval Informativeness.
+CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
+Eccen.
+EigVal.
+LexSim.
+NumSet.
+1S 2S 4-shot 8-shot 16-shot Judge(Func.) 0.40 0.36 0.31 0.36 0.36 0.17 0.20 0.29 0.31 0.33 Judge(Incon.) 0.46 0.31 0.29 0.34 0.26 0.13 0.17 -0.14 -0.14 -0.10 APIGen CodeBLEU 0.59 0.33 0.38 0.41 0.10 -0.01 0.15 0.08 0.05 0.06 ASQA LLMEval(A.) 0.33 0.24 0.35 0.36 0.30 0.33 0.15 0.24 0.22 0.25 LLMEval(I.) 0.39 0.27 0.38 0.41 0.30 0.28 0.15 0.18 0.28 0.14 ASQA BERTScore 0.51 0.34 0.48 0.57 0.10 0.33 0.23 0.04 0.07 0.11 BiGGen Judge 0.44 0.13 0.44 0.50 0.27 0.24 0.21 0.19 0.08 0.16 BigCodeBench Judge(Func.) 0.34 0.32 0.33 0.33 0.05 0.10 0.04 0.07 0.10 0.08 Judge(Incon.) 0.36 0.35 0.35 0.35 0.10 0.06 -0.01 -0.08 -0.08 -0.08 BigCodeBench CodeBLEU 0.58 0.50 0.56 0.60 0.12 0.05 0.02 0.04 0.09 0.05 LLMEval(A.) 0.48 0.19 0.49 0.50 0.27 0.12 0.21 0.13 0.04 0.10 LLMEval(I.) 0.46 0.20 0.50 0.50 0.26 0.10 0.18 0.19 0.14 0.18 Eli5 BERTScore 0.58 0.27 0.57 0.59 -0.01 -0.03 0.08 0.04 0.06 0.07 Fact-Bio FActScore 0.78 0.31 0.78 0.75 0.30 0.65 0.21 0.45 0.56 0.54 LLMEval(A.) 0.42 0.17 0.43 0.45 0.23 0.23 0.13 0.10 0.08 0.13 LLMEval(I.) 0.42 0.16 0.39 0.43 0.13 0.24 0.16 0.20 0.22 0.24 MED LF QA BERTScore 0.55 0.38 0.52 0.54 0.16 0.22 0.36 0.09 0.04 0.05 Opus-100 GEMBA 0.50 0.36 0.44 0.49 0.28 0.33 0.21 0.16 0.21 0.22 Opus-100 BERTScore 0.66 0.50 0.63 0.65 0.23 0.11 0.15 0.06 0.05 -0.02 G-Eval(Coh.) 0.01 0.05 0.03 0.02 0.16 0.05 0.04 0.06 -0.01 -0.02 G-Eval(Con.) 0.01 0.09 0.10 0.05 0.15 0.08 0.02 0.02 0.03 -0.01 G-Eval(Flu.) 0.09 0.01 0.10 0.05 0.03 0.03 0.03 0.00 -0.02 0.02 G-Eval(Rel.) 0.15 0.09 0.17 0.13 0.14 0.18 0.05 0.04 0.01 0.06 Summeval BERTScore 0.56 0.24 0.57 0.52 0.08 0.16 0.00 -0.03 0.04 -0.02 LLMEval(A.) 0.53 0.27 0.55 0.55 0.08 0.25 0.22 0.04 0.08 0.06 LLMEval(I.) 0.48 0.24 0.45 0.49 0.06 0.14 0.15 0.07 0.04 0.12 TruthfulQA BERTScore 0.60 0.13 0.39 0.46 0.08 0.12 0.11 0.03 0.10 -0.02 AVERAGE 0.43 0.25 0.41 0.42 0.17 0.17 0.13 0.09 0.10 0.08
+Table 9: Pearson Correlation Benchmark.
+We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16-shot settings.
+For RMSE, lower is better.
+Judge(Func.) = CodeJudge Functional Correctness.
+Judge(Incon.) = CodeJudge Inconsistency.
+LLMEval(A.) = LLM Eval Accuracy.
+LLMEval(I.) = LLMEval Informativeness.
 CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
@@
 LLMEval(I.) = LLMEval Informativeness.
-CE-Reg: Consistency-based CE-Reg: Verbalized RF-LLMaaJ Dataset Metric DegMat.
-Eccen.
-EigVal.
-LexSim.
-NumSet.
-1S 2S 4-shot 8-shots 16-shot APIGen Judge(Func.) 0.48 0.38 0.38 0.42 0.34 0.19 0.16 0.25 0.27 0.23 Judge(Incon.) 0.55 0.35 0.30 0.35 0.17 0.09 0.13 0.03 0.01 0.08 CodeBLEU 0.65 0.38 0.41 0.47 0.08 0.02 0.11 0.05 0.04 0.01 ASQA LLMEval(A.) 0.47 0.30 0.43 0.52 0.23 0.41 0.35 0.14 0.14 0.11 LLMEval(I.) 0.38 0.27 0.36 0.42 0.16 0.33 0.27 0.02 0.05 0.08 BERTScore 0.51 0.35 0.53 0.58 0.07 0.34 0.16 0.07 0.05 0.10 BiGGen Judge 0.62 0.26 0.52 0.55 0.10 0.21 0.25 0.26 0.21 0.06 BigCodeBench Judge(Func.) 0.46 0.36 0.41 0.48 0.07 0.11 0.15 0.06 0.01 0.02 Judge(Incon.) 0.33 0.27 0.26 0.35 0.04 0.08 0.13 0.09 0.02 0.04 CodeBLEU 0.56 0.47 0.50 0.56 0.09 0.00 0.05 0.07 0.03 -0.07 Eli5 LLMEval(A.) 0.50 0.18 0.48 0.46 0.31 0.16 0.21 0.12 0.09 0.15 LLMEval(I.) 0.36 0.13 0.31 0.30 0.23 0.17 0.22 0.22 0.24 0.18 BERTScore 0.54 0.32 0.53 0.51 0.06 0.01 0.11 -0.02 -0.04 0.02 Fact-Bio FActScore 0.72 0.25 0.74 0.77 0.27 0.68 0.26 0.47 0.50 0.54 MED LF QA LLMEval(A.) 0.48 0.24 0.49 0.50 0.24 0.31 0.24 0.11 0.13 0.13 LLMEval(I.) 0.44 0.30 0.42 0.43 0.07 0.28 0.24 0.24 0.28 0.27 BERTScore 0.55 0.38 0.52 0.54 0.16 0.22 0.36 0.09 0.04 0.05 Opus-100 GEMBA 0.52 0.37 0.48 0.55 0.29 0.24 0.14 0.09 0.11 0.08 BERTScore 0.66 0.50 0.63 0.65 0.23 0.11 0.15 0.06 0.05 -0.02 Summeval G-Eval(Coh.) 0.38 0.17 0.36 0.35 0.09 0.23 0.02 0.00 0.04 0.02 G-Eval(Con.) 0.29 0.10 0.34 0.30 0.12 0.17 0.02 -0.02 -0.04 -0.07 G-Eval(Flu.) 0.08 0.00 0.08 0.08 0.08 0.04 -0.03 -0.04 0.05 -0.03 G-Eval(Rel.) 0.40 0.13 0.38 0.38 0.16 0.23 0.02 0.02 0.03 0.01 BERTScore 0.56 0.24 0.57 0.52 0.08 0.16 0.00 -0.03 0.04 -0.02 TruthfulQA LLMEval(A.) 0.60 0.24 0.54 0.54 0.12 0.18 0.31 0.06 -0.01 0.12 LLMEval(I.) 0.43 0.21 0.35 0.39 0.14 0.20 0.18 0.15 0.06 0.20 BERTScore 0.60 0.13 0.39 0.46 0.08 0.12 0.11 0.03 0.10 -0.02 AVERAGE 0.49 0.27 0.43 0.46 0.15 0.20 0.16 0.10 0.09 0.08
-Table 9: PC Benchmark.
-We compare two baselines: 1) confidence-based regression (CE-Reg) uses 7 CE features (five consistency-based, two verbalized); and 2) reference-free LLM-as-a-Judge (RF-LLMaaJ), evaluated in 4-, 8-, and 16- settings.
-For PC, values higher the better.
-Judge(Func.) = CodeJudge Functional Correctness.
-Judge(Incon.) = CodeJudge Inconsistency.
-LLMEval(A.) = LLM Eval Accuracy.
-LLMEval(I.) = LLMEval Informativeness.